### v1

# revtex4-2Repair the float

Abstract
Lying at the interface between Network Science and Machine Learning, node embedding algorithms take a graph as input and encode its structure onto output vectors that represent nodes in an abstract geometric space, enabling various vector-based downstream tasks such as network modelling, visualization, data compression, node classification, link prediction, and community detection. Two apparently unrelated limitations affect these algorithms. On one hand, it is not clear what the basic operation defining vector spaces, i.e. the vector sum, corresponds to in terms of the original nodes in the network. On the other hand, while the same input network can be represented at multiple levels of resolution by coarse-graining the constituent nodes into arbitrary block-nodes, the relationship between node embeddings obtained at different hierarchical levels is not understood. Here, building on recent results in network renormalization theory, we address these two limitations at once and define a multiscale node embedding method that, upon arbitrary coarse-grainings, ensures statistical consistency of the embedding vector of a block-node with the sum of the embedding vectors of its constituent nodes. We illustrate the power of this approach on two economic networks that can be naturally represented at multiple resolution levels: namely, the network of international trade between (sets of) countries and the network of input-output flows among (sets of) industries in the Netherlands. We confirm the statistical consistency between networks retrieved from coarse-grained node vectors and networks retrieved from sums of fine-grained node vectors, a result that cannot be achieved by alternative methods. Several key network properties, including a large number of triangles, are successfully replicated already from embeddings of very low dimensionality, allowing for the generation of faithful replicas of the original networks at arbitrary resolution levels.
I Introduction
Complex networks capture a variety of socially relevant processes, from economic activities to interactions among brain regions [1, 2, 3]. Indeed, every dyadic interaction can be described by properly defining which are the actors (nodes) and the type of connections among them (edges). By defining the nodes as the sectors and the edges as the transactions among them, the Input-Output network (ION) [4] is recovered. Again, by defining the nodes as the states and the edges as the trade among them, the World Trade Web (WTW) [5] is obtained. On the other hand, if nodes were seen as the brain regions connected by electrical stimuli, the brain network is represented. In particular, this flexibility allows for arbitrary definitions of the nodes even when looking at the same phenomenon that is generating the data (generative process). For example, one could have access to the network involving the most detailed sector classification (National Industry), whereas another one only at the community level (Industry) - see the upper part of Figure 1. The same reasoning applies naturally in other contexts (neuroscience, social sciences, $\cdots$), as community structures help in simplifying the heterogeneity of the graph [6, 7]. Therefore, seen at a coarser resolution, the graph represents the interactions among block-nodes, and it would be uniquely recovered after the specification of the partitions. This scheme could be iterated at wish to produce a multi-scale unfolding of the original graph with nested partitions: pictorially, the base of a pyramid is the observed network, whereas the coarser levels are the cross-sections of the pyramid. Lastly, it is worth noticing that the properties of each lower-resolution graph change with levels. For instance, the firm graph is less dense than the sector one since there will be fewer nodes to redistribute links to.
Here, assuming to know the nodes, we aim at modeling their interactions by assigning a probability for every pair of nodes (or edge or link) [8]: the higher the probability, the more likely is that edge to exist in a sampled graph. Ultimately, the measurements over the observed graph should coincide with the average over the sampled networks. This exercise is called, in general, network modeling but also network reconstruction [9] or (binary) edge classification [10] in the machine-learning literature.
To tackle this problem, many machine-learning models use node embeddings [11, 12, 2]. As highlighted by the arrows in Figure 1, one may assign a set of coordinates for every node and, then, extract the probability for the network’s edges. The optimal node embeddings are the vector of parameters that optimize a functional involving the observed graph and the model, for e.g. the likelihood. Therefore, these vectors are interpreted as informative features about the nodes that can be deployed to different tasks, such as community detection or node classification[13].
Two hallmarks of real-world graphs are 1) low density (sparsity) and 2) high triangle density, where there are many triangles incident to low-degree vertices [14]. It was recently proven that linear node embedding methods, such as node2vec [11], are not capable of reproducing the triangle density [14]. To overcome this drawback, LogisticPCA [15] optimized the vectors for a non-linear logistic function. Furthermore, Chanpuriya et colleagues proposed the “symmetric” LogisticPCA (LPCA) [16] to deal with undirected networks (see subsection III.1).
As mentioned, a phenomenon can be studied at different scales. By combining nodes into communities we go from a microscopic to a coarse grained scale. Nevertheless, most of the models, e.g. LPCA, regard a network only at one scale, providing the optimal embeddings for that level. If nodes were merged into communities, the block-vectors have to be recomputed. In other words, the two sets of vectors are completely unrelated, as if the models see the two networks as realizations of two distinct processes - even though it is not the case. there is no prescription to use the node embeddings from a lower level to create an embedding for a community. For this reason, we would refer to this class of models as “single-scale” models (SSM).
To tackle the renormalization problem on networks, several methods have been proposed in the literature, but they all rely on strong assumptions that limit their use cases [17]. More concretely, the most promising one [3] assumes that the nodes are embedded in a hyperbolic plane, all the coarser networks are scale-free and the block-nodes contain $r=2$ micro-nodes. The latter restriction doesn’t allow to aggregate the nodes with the “most natural” way induced by the studied phenomenon, e.g. geographical distances for the WTW or the sector (industry) for the ION. To overcome this limitation, the multi-scale model [17] was proposed, which is generalizable at higher levels and allows for any arbitrary partition of the microscopic nodes. Indeed, the block parameters are uniquely obtained by summing the vectors of the nodes belonging to a community (renormalization rule) - this mimics the unique identification of the coarser network starting from the microscopic one.
In this work, we enhance the scalar multi-scale model with node embeddings (MSM). The renormalization rule for vectors is shown in Figure 1: each community embedding is the summation of its lower-level ones. Therefore, the MSM provides an interpretation of the sum of node embeddings, which is rarely addressed in the literature [^1] The successes of Natural Language Processing are also due to the effective representation of a phrase obtained by summing the embedding for each word of it [18]. This is hardly replicated in a graph setting as it is intrinsically more complex than a language. [19]. Due to the renormalization rule, the MSM has the additional benefit of having to be fitted only at the ground level, further implying a lower computational complexity with respect to a single-scale model to be tailored at every level.
The rest of the paper is organized as follows. In section III, we introduce the LogisticPCA and the Multi-Scale Model (MSM) alongside its renormalization rule. In subsection IV.1, we describe the ING Input-Output Network and the World Trade Web datasets. Furthermore, we present the coarse-graining procedure to obtain the higher-scale network. In section VI, we show the multi-scale results of the two models and discuss the implications of the theoretical results over either network- and machine-learning scores. Finally, in the Appendices, we store all the technical details supporting the results in the main text.
Figure 1: Visualization of the research question: how node embeddings relate across scales? Here, it is shown the embedding extraction scheme for two levels: on the left the microscopic level that uniquely generates the coarser one on the right after aggregation. However, for SSM, the vectors fitted at the microscopic level can’t be used to obtain the ones computed at level 1 (red question marks), whereas its the case for our MSM. The latter succes is possible thanks to the vector sum which is highlighted by using the parallelogram law on the right, e.g. pale red vectors are aligned in order to get the solid red vector at an higher scale.
II Graph Renormalization
In this section, we will introduce the mathematical description of a graph and its coarse-graining procedure. We use the subscript $\ell=0$ to identify the base graph whereas $\ell>0$ for its aggregated versions. In addition, we refer to a generic quantity at level $\ell$ as $\ell$-quantity, e.g. the 2-vectors and 2-nodes are, respectively, the node embeddings and block-nodes at level $2$.
Consider a binary undirected graph with $N_{0}$ “microscopic” nodes $\mathcal{V}_{0}$ (labelled as $i_{0}\in[1,N_{0}]$) and their connections (called edges or links), i.e.
$\mathcal{E}_{0}:=\left\{(i_{0},j_{0}):i_{0}\in\mathcal{V}_{0},j_{0}\in[i_{0},N_{0}],a^{(0)}_{i_{0}j_{0}}=1\right\}$
where $a^{(0)}_{i_{0}j_{0}}=1$ if there was an edge among $i_{0},j_{0}$ and $a^{(0)}_{i_{0}j_{0}}=1$ otherwise. This system could be represented by an $N_{0}\times N_{0}$ adjacency matrix $\mathbf{A}^{(0)}$ which has to be symmetric, i.e. $a^{(0)}_{i_{0}j_{0}}=a^{(0)}_{j_{0}i_{0}}$, since the graph is undirected. We don’t account for multiple edges, whereas we do for self-loops in the diagonal of $\mathbf{A}^{(0)}$, i.e. $a^{(0)}_{i_{0}i_{0}}=0,1$. Unless explicitely specified, we would use the notation $\xi_{ij}:=\xi^{(\ell)}_{i_{\ell}j_{\ell}}$ with $\xi$ as a chosen quantity when the results are valid for any level $\ell\geq 0$, e.g. if $\xi:=\mathbf{A}^{(\ell)}\textnormal{ then }a_{ij}:=a^{(\ell)}_{i_{\ell}j_{\ell}}$.
II.1 Coarse-Graining
In order find the coarse-grained version of the graph, we define non-overlapping arbitrary partitions $\Omega_{0}$ of the microscopic nodes into block-nodes
$\mathcal{V}_{1}:=\left\{i_{1}:=\Omega_{0}(i_{0})\quad\forall i_{0}\in\mathcal{V}_{0}\right\}.$
Secondly, we assume that two communities are connected if there was at least one edge between their internal nodes, i.e.
$a^{(1)}_{i_{1}j_{1}}=1-\prod_{i_{0}\in\Omega_{0}^{-1}(i_{1}),j_{0}\in\Omega_{0}^{-1}(j_{1})}(1-a^{(0)}_{i_{0}j_{0}})$ (1)
where the operation used is the logical OR over the edges connecting two communities. That is, only the zeros are preserved whereas the possible multiple-edges among the lumped nodes are projected to one. Moreover, note that $\Omega_{0}$ is only surjective but not injective, since multiple nodes $i_{0}$ are merged into $i_{1}$, and that we allow for self-loops as we didn’t require $i_{1}\neq j_{1}$. At a higher level, the self-loop proxies there is at least one connection inside the community either because links among nodes or self-loops edges. In this way, we built the $N_{1}\times N_{1}$ adjacency matrix $\mathbf{A}^{(1)}$ which is binary and symmetric as $\mathbf{A}^{(0)}$. To ease the recalling, we will refer to $A^{(0)}$ as 0-graph, $A^{(1)}$ as 1-graph, the $N_{0}$ microscopic nodes as $0-nodes$ and the block-nodes as $1-nodes$.
The lumping procedure could be iterated $\ell+1(\geq 1)$ times by specifying the partition $\Omega_{\ell}$ of $\ell-nodes$ into $N_{\ell+1}$ block-nodes
$\mathcal{V}_{\ell+1}:=\left\{i_{\ell+1}:=\Omega_{\ell}(i_{\ell})\,\forall i_{\ell}\in\mathcal{V}_{\ell}\right\}.$
However, since the partitions $\left\{\Omega_{\ell}\right\}_{\ell\geq 0}$ are not overlapping, one can define their composition $\Omega_{0\to\ell}$ as
$\Omega_{0\to\ell}:=\Omega_{\ell}\circ\,\cdots\,\circ\Omega_{0}.$ (2)
This provides a direct mapping of $0-nodes$ into block-nodes $i_{\ell+1}=\Omega_{0\to\ell}(i_{0})$. Therefore, the $(\ell+1)-graph\textnormal{, }\mathbf{A}^{(\ell+1)}$ is recovered both applying iteratively the Equation 1 or directly via Equation 2
$a^{(\ell+1)}_{i_{\ell+1}j_{\ell+1}}$ $=1-\prod_{i_{\ell}\in\Omega_{\ell}^{-1}(i_{\ell+1}),j_{\ell}\in\Omega_{\ell}^{-1}(j_{\ell+1})}(1-a^{(\ell)}_{i_{\ell}j_{\ell}})$ (3)
$=1-\prod_{i_{0}\in\Omega^{-1}_{\ell\to 0}(i_{\ell+1}),j_{\ell}\in\Omega^{-1}_{\ell\to 0}(j_{\ell+1})}(1-a^{(0)}_{i_{0}j_{0}})$ (4)
where $\Omega^{-1}_{\ell\to 0}:=\Omega^{-1}_{0}\circ\dots\circ\Omega^{-1}_{\ell}$ is the inverse of Equation 2.
The nested $\left\{\Omega_{\ell}\right\}_{\ell\geq 0}$ can be uniquely parametrized in terms of a dendrogram as shown in [17]. By following an “horizontal” cut of the dendrogram, one obtains the community partitions at the same scale, e.g. for the WTW by merging states nearer than a fixed geographical distance. In contrast, by cutting the dendrogram at different heights, one recovers the “multiscale” clusters, e.g. a state interacting with a continent.
In many cases, the problem setting suggests a partitioning of the nodes into communities without an explicit distance matrix, e.g. the classification of the Industries based on the NAICS-codes digits. They characterize each Industry by a number at $6$ digits, e.g. Full-Service Restaurants (722511), but also their clusters by progressively removing the last number up to $2$ digits, e.g. Accommodation and Food Services (72). More concretely, the industries sharing the first $n\in[2,5]$ digits are merged into a block-node at the $\ell:=6-n\in[1,4]$ level (cfr. Hamming Distance). The dendrogram can be built by setting the height where the blocks split into sub-cluster as $\ell$.
Lastly, if the problem doesn’t come with any “natural” partition, it would be possible to merge the 0-nodes randomly by creating a dendrogram from a random distance matrix.
III Models
In the previous sections, we highlighted that one phenomenon could be studied at different resolutions. In particular, since every adjacency matrix $\mathbf{A}^{(\ell)}$ is binary and undirected, every edge $(i_{\ell},j_{\ell})$ could be seen as a Bernoulli random variable. Since the theory is valid for every scale, we will remove the dependence on the level $\ell$ on each quantity, e.g. $a_{ij}:=a^{(\ell)}_{i_{\ell}j_{\ell}}$ (see section II). Hence, for every resolution the adjacency entries are
$a_{ij}=\begin{cases}1&\qquad p_{ij}\\ 0&\qquad 1-p_{ij}\end{cases}\qquad\forall i\leq j.$ (5)
To model this hierarchical structure, we will use the LogisticPCA [16] and the new MSM. We selected LPCA as the representative for the SSM whereas, at the best of our knowledge, only the MSM accounts for a renormalization rule over arbitrary partitions of the nodes.
III.1 Logistic PCA
LPCA [^2] The authors have historically called LPCA the directed LPCA, but there was no straightforward acronym for its undirected counterpart [16]. Since in this work, we will use only the symmetric one, we will call it LPCA for easiness. [16] is a probabilistic model which aims to classify each edge as existing (class $1$) or non-existing (class 00). This is a common “spam / non-spam” problem [13], but applied to objects (the links) that are not carrying explicit features - that is why node embeddings come into play.
The connection probability reads
$p_{ij}:$ $=\sigma(\vec{b}_{i},\vec{b}_{j},\vec{c}_{i},\vec{c}_{j})$ (6)
$=\frac{1}{1+e^{-\left(\langle\vec{b}_{i},\vec{b}_{j}\rangle-\langle\vec{c}_{i},\vec{c}_{j}\rangle\right)}}$ (7)
where the node embeddings
$\vec{b}_{i}\in\mathbb{R}_{+}^{D_{B}\geq 1},\vec{c}_{i}\in\mathbb{R}_{+}^{D_{C}\geq 1}$
were introduced by the Chanpuriya et colleagues to grasp the “homophily” and “heterophily” natures of a node respectively (see Supp.Mat. I). By defining the non-negative matrices as $\textbf{B}:=\begin{bmatrix}-\vec{b}_{1}^{T}-\\ \vdots\\ -\vec{b}_{N_{0}}^{T}-\end{bmatrix}$ and similarly for $\mathbf{C}$, the scalar product among them becomes $\mathbf{X}:=\mathbf{B}\mathbf{B}^{T}-\mathbf{C}\mathbf{C}^{T}$ implying $x_{ij}:=\langle\vec{b}_{i},\vec{b}_{j}\rangle-\langle\vec{c}_{i},\vec{c}_{j}\rangle\in\mathbb{R}$. Therefore, the node embeddings are obtained by maximizing the likelihood [20]
$\mathcal{L}\left(\mathbf{X}|\mathbf{A}\right)$ $:=\mathcal{L}\left(\mathbf{B},\mathbf{C}|\mathbf{A}\right)=$ (8)
$=\sum_{i\leq j}\min(x_{ij},0)$ $-\ln(1+e^{-\left|x_{ij}\right|})-(1-a_{ij})x_{ij}$ (9)
subject to $\vec{b}_{ik}\geq 0,\vec{c}_{ik}\geq 0$.
III.2 Multi-Scale Model
The natural way of modelling $\left\{\mathbf{A}^{(\ell)}:\ell\geq 0\right\}$ is deploying renormalization models that have to be self-consistent across scales and allowing for arbitrary partitions $\left\{\Omega_{\ell}\right\}_{\ell\geq 0}$. The multi-scale model [17] is the only one that takes into account all these properties (cfr. [3] for homogeneous partitions). In addition, by equipping the (global) multi-scale model with node embedding $\left\{\vec{x}_{i}\right\}_{i\in[1,N]}$, the connection probability for the new multi-scale model (MSM) reads
$p_{ij}:=p(\vec{x}_{i},\vec{x}_{j},w_{i})=\begin{cases}\makebox[51.96254pt][l]{$1-e^{-\langle\vec{x}_{i},\vec{x}_{j}\rangle}$}i\neq j\\ \makebox[51.96254pt][l]{$1-e^{-\frac{1}{2}\|\vec{x}_{i}\|^{2}-w_{i}}$}i=j\end{cases}$ (10)
subject to $\vec{x}_{ik}\geq 0,w_{i}\geq-\frac{1}{2}\|\vec{x}_{i}\|^{2}$ since the probability has to be bounded ($p_{ij}\in[0,1]$). In principle, one should impose $\langle\vec{x}_{i},\vec{x}_{j}\rangle\geq 0$, but this is equivalent to restricting every component to be non-negative, i.e. $\vec{x}_{ik}\geq 0$ (see Supp.Mat. II.6). The norm $\|\vec{x}_{i}\|:=\langle\vec{x}_{i},\vec{x}_{i}\rangle$ is the one induced by the scalar product. In general, $\vec{x}_{i}$ could be interpret as the propensity of $i$ to create a link with another node $j$. Theoretically, this role in encoded by its modulus (significance) and its relative angle (similarity) with the other embeddings. On the other hand, $w_{i}$ rules the self-loop of node $i$. A more detailed interpretantion of the node embeddings is lacking in the literature, but it is out of the scope for this work.
As for LPCA, $\left\{\vec{x}_{i}\right\}_{[0,N-1]}$ are obtained by maximizing the likelihood
$\mathcal{L}\left(\mathbf{X}|\mathbf{A}\right)$ $:=\sum_{i\leq j}a_{ij}\ln\left(p_{ij}\right)+\left(1-a_{ij}\right)\ln\left(1-p_{ij}\right)$ (11)
$=\sum_{i\leq j}a_{ij}\ln\left(1-e^{-\langle\vec{x}_{i},\vec{x}_{j}\rangle}\right)-\left(1-a_{ij}\right)\langle\vec{x}_{i},\vec{x}_{j}\rangle$ (12)
where $\textbf{X}:=\begin{bmatrix}-\vec{x}_{1}^{T}-\\ \vdots\\ -\vec{x}_{N}^{T}-\end{bmatrix}\in\mathbb{R}^{N\times D}$ is the vertical stacking of the embedding vectors, i.e. $\mathbf{X}_{i*}=\vec{x}_{i}$. We optimized only the vectors of the structural inequivalent nodes (see Supp.Mat. II.4). The loop parameters $\left\{w_{i}\right\}_{i\in[1,N]}$ are fixed to
$w_{i}=\begin{cases}\makebox[28.10083pt][l]{$-\frac{1}{2}\|\vec{x}_{i}\|^{2}$}\quad\textnormal{ if }p_{ii}=1\\ \makebox[28.10083pt][l]{$\to\infty$}\quad\textnormal{ if }p_{ii}=0\end{cases}$ (13)
after the training of the node embeddings. Indeed, for each node $i$, the likelihood of the self-loops $i$ displays only one term related to the presence ($a_{ii}=1$) or absence of the self-loop ($a_{ii}=0$). Therefore, by maximizing the likelihood the reached values for the loop parameters will be exactly the ones set above.
III.2.1 Self-Consistency and Renormalization
The MSM enforces the equality among the renormalized $\widetilde{\mathbf{P}}^{(\ell)}$ and the coarse-grained $\check{\mathbf{P}}^{(\ell)}$ probabilities [^3] The notation for the different probabilities uses the $\hat{}$ (hat) to refer to the P with fitted parameters. On the contrary, the $\;\check{}\;$ (check) reuses the 0-level parameters (the opposite of $\hat{\quad}$), and the $\widetilde{\quad}$ (tilde) resembles the “S” of “summed” for the renormalized probabilities. with the same functional form ([17], Supp.Mat. II), namely
$\begin{cases}\widetilde{\mathbf{P}}^{(\ell)}\stackrel{{\scriptstyle!}}{{=}}\check{\mathbf{P}}^{(\ell)}\\[4.30554pt] \mathbf{P}^{(\ell)}\stackrel{{\scriptstyle!}}{{=}}\mathbf{P}^{(m)}\qquad\forall\ell\geq 0,m\geq 0\end{cases}$ (14)
These two conditions are called self-consistency and scale-invariance. Here, we will use again the subscript $\ell=0$ to highlight that $\left\{\vec{x}_{i_{0}},w_{i_{0}}\right\}_{i\in[1,N_{0}]}$ are the fitted vectors at finest level 00, whereas $\left\{\vec{x}_{I},w_{I}\right\}_{I\in[1,N_{\ell}]}$ are the parameters at the level $\ell\geq 0$ where we set the block-nodes as $I:=\Omega_{0\to\ell}(i_{0})$. Therefore, by assuming that the block-node parameters are the sum of the lower level ones, i.e.
$\vec{x}_{I}=\sum_{i_{0}\in I}\vec{x}_{i_{0}}\textnormal{ and }w_{I}=\sum_{i_{0}\in I}w_{i_{0}},$ (15)
Equation 14 are satisfied by means of Equation 10 (see section II). In this way, the MSM at level $\ell$ is recovered by inserting them in the Equation 10, namely
$p_{IJ}:=p(\vec{x}_{I},\vec{x}_{J},w_{I})=\begin{cases}\makebox[45.8509pt][l]{$1-e^{-\langle\vec{x}_{I},\vec{x}_{J}\rangle}$}I\neq J\\ \makebox[45.8509pt][l]{$1-e^{-\frac{1}{2}\|\vec{x}_{I}\|^{2}-w_{I}}$}I=J\end{cases}$ (16)
Thanks to the summation rule, the MSM has a lower computational complexity with respect to the single-scale models (see subsection II.7) as they need to be refitted at every scale.
In this essay, we took advantage of this property by fitting the embeddings vector from the observed network to recover all the coarser $\ell-parameters$ for every $\ell\geq 0$ by means of Equation 15. Finally, the probability $p_{IJ}$ among $\ell-communities$ was obtained by inserting the $\left\{\vec{x}_{I},w_{I}\right\}_{I\in[1,N_{\ell}]}$ in Equation 10.
IV Applications
IV.1 ING Input-Output Network
ING Bank N.V. regularly reports the economic transactions of all ING clients for different years. We focused on the payments for the year 2022 between ING firms by removing the individual or non-Dutch clients, the flows of money sent/received by a non-ING account and the payments circulating inside a firm, i.e. self-payments. Since ING is the biggest bank in The Netherlands [21], this gave us the possibility of analyzing a major portion of the market.
More precisely, we chose the year 2022 both to ease the numerical calculations and to avoid skewed distributions by the aftermath of the COVID-19 pandemic. However, the procedure may be easily replicated for other time intervals, e.g. 3 years span, quarterly, $\cdots$
At the firm-to-firm (f2f) resolution, the dataset is composed by $N_{f2f}\approx 3.4\cdot 10^{5}\text{ nodes},L_{f2f}\approx 4\cdot 10^{4}\text{links}$ which imply a density $\rho_{f2f}\approx 3.5\cdot 10^{-5}$. Therefore, the network is big and sparse. From this network, we aggregated the firms by NAICS (North American Industry Classification System) codes and set an edge among two sectors if there was at least one link among the firms of each community (see Equation 3). Then, we filtered out the “Public Administration” (92), “Finance and Insurance” (52), “Management of Companies and Enterprises”/“Holdings” (55) sectors to retain a production ION. Roughly, the reasons underneath the payments from/to the sectors 52-55-92 are not directly connected with a product/service. In particular, the “Public Administration” fluxes includes taxes and fees; the “Finance and Insurance”, money management, e.g. loans, that are not part of the production chain of any good; and the “Management of Companies and Enterprises”/“Holdings” are a collection of business entities, controlling stocks in other companies. Lastly, we mapped, for simplicity, the 6-digits NAICS codes to integers, i.e. $111110\to 0,111120\to 1,\cdots$.
This is a first application of the MSM on the multiscale structure built from the ION. Therefore, we studied only the economic relationships among the sectors discarding the directionality and the amount of money of the link. In other words, each edge is reciprocated and binary. For example [^4] To be precise, the notation for the nodes $i,j$ refers to the observed microscopic nodes, namely $i:=i_{0},j:=j_{0}$, if $w_{ij}$ was the total amount of money sent from $i\textnormal{ to }j$ at level 00, we reciprocated the weight by setting $w_{ij}\to w^{{}^{\prime}}_{ij}:=\frac{w_{ij}+w_{ji}}{2}$ [5] and $a_{ij}=a_{ji}=1$ if $w^{{}^{\prime}}_{ij}>0$ and imposed $a_{ij}=a_{ji}=0\textnormal{ if }w^{{}^{\prime}}_{ij}=0$ to create a binary edge. In this way, a bidirectional link is created, i.e. $a_{ij}=1$ every time there was at least one directed flow among two sectors. The empirical ION is composed by $N_{0}=972$ sectors and $L_{0}=1.4\cdot 10^{5}$ links, which imply a density $\rho_{0}\approx 0.29\gg\rho_{f2f}$ by 4 order of magnitudes.
IV.1.1 Coarse-Graining The ION
In the previous section, we obtained the binary undirected adjacency matrix $\mathbf{A}^{(\ell=0)}$ representing the interactions among the 6-digits sectors (0-nodes). Here, we describe the coarse-graining procedure producing the multi-scale unfolding of $\mathbf{A}^{(0)}$.
[…]
Secondly, we set an edge among two $\ell-nodes$ if there was at least one link among their higher resolution members. More formally, the coarse-grained $\mathbf{A}^{(\ell)}$ is calculated by applying the Equation 4. We chose this one-step procedure to easily generate every level without passing into the intermediate scales. However, the MSM accept every arbitrary-steps scheme, e.g. if $\ell=3$ then $0\to 1\to 3$ is also possible. Note that the coarse-grained graphs become fully-connected from $\ell=4$. Therefore, the statistical modelling would be possible up to $\ell=3$.
IV.2 World Trade Web
As a second application, we considered the World Trade Web (WTW) from the Gleditsch dataset [23], which reports the international trade flows (imports and exports) among all the world countries. We selected the year 2000 (the most recent one) and removed the states that were not reported in the BACI-CEPII GeoDist [24] as we would use the geographical distances to coarse-grain the WTW. This results in $N_{0}=185$ 0-nodes. Although we analyzed the year 2000 the methodology can be applied also to other years.
The dataset provides two columns which display, for every pair of nodes $(i,j)$: the export $w_{ij}$ and import $w_{ji}$ among $i,j$ in USA dollars. However, by flipping $i\leftrightarrow j$, the export $\omega_{ji}$ and import $\omega_{ij}$ don’t coincide with the previous value. Therefore, we used only the redefined $w_{ij}\leftarrow\frac{w_{ij}+\omega_{ji}}{2}$ [25] as the amount of trade from $i\textnormal{ to }j$.
As done for the ION, we symmetrized the connections by mapping $w_{ij}\to w^{{}^{\prime}}_{ij}:=\frac{1}{2}\left(w_{ij}+w_{ji}\right)$, i.e. the average flow between the two directions. By renaming $w^{{}^{\prime}}_{ij}$ as $w_{ij}$, it follows that the weights are symmetric, i.e. $w_{ij}=w_{ji}$. Note that the WTW has a high reciprocity of links, i.e. the connections in the WTW are almost always corresponded [26]. Therefore, its undirected approximation is a legitimate starting point to study the system. Moreover, we binarized the import-export matrix to get an adjacency matrix among the states. Formally, we projected all the positive values of the weighted matrix $\textbf{W}:=\left\{w_{ij}\right\}_{i\in[1,N_{0}]}$ to one, i.e. $\mathbf{A}:=\Theta(\textbf{W})$ where the $\Theta(\cdot)$ is the Heaviside function.
IV.2.1 Coarse-Graining of WTW
To coarse-grain the empirical WTW, assumed at level 00, we used the geographical distances [24] to iteratively merge “close” nodes into block-nodes as in [17]. Technically, this is done by means of a single-linkage agglomerative clustering model which returns a dendrogram where the leaves are the 0-nodes, the branching points are the block-countries, and the height of each branching point represents the distance, obtained via the single-linkage, between two leaves following the corresponding branches. Therefore, to calculate the partitions $\left\{\Omega_{\ell}\right\}_{\ell\geq 0}$ we cut the dendrogram at $18$ hierarchical heights $\ell\in[0,17]$, such that the number of block-countries, namely $i_{\ell+1}:=\Omega_{\ell}(i_{\ell})$, are $N_{\ell}:=N_{0}-10\cdot\ell$. Note that the coarse-grained graphs become fully-connected from $\ell=7$. Therefore, the statistical modelling would be possible up to $\ell=6$.
V Methodology
In this section, we will define the scores used to evaluate the models over the ION and the WTW.
V.1 Embedding Dimension
The choice of the best embedding dimension is still an active Research topic [27, 28, 29]. Here, we relied on the “Minimum Description Length” principle approximated by the Bayesian Information Criteria (BIC) [28]. In the Supp.Mat. Table 1, we reported the BIC scores over the dimensions $D=1,2,8,16$ and also, for completeness, the AIC ones [27]. The best dimension, according to BIC, depends on the resolution levels: for the ION, $D^{best}_{\ell\geq 1}=1$ whereas $D^{best}_{\ell=0}=2$. Since the WTW is a less complex network, the BIC selects everytime the lowest dimension, i.e. $D^{best}_{\ell\geq 0}=1$ (see Supp. Mat.). This result could be expected since having more nodes, as for the lower levels, it implies more heterogeneity and, therefore, a bigger embedding dimension to grasp the ION. However, since this is the first time the MSM has been proposed, we will show the scores for all the $D=1,2,8,16$ dimensions.
V.2 Renormalization of LPCA
Having fitted the $\vec{b}_{i_{0}},\vec{c}_{i_{0}}$ (at level 00), in order to model $A^{(\ell)}$, one may either proceed with the Equation 8 or by refitting LPCA at that level. However, both solutions have similar drawbacks as there is no renormalization rule enforcing the scale-invariance (cfr. Equation 15). In the following, we will analyze them in detail.
Firstly, one may calculate $\check{\mathbf{P}}^{(\ell)}$ from the RHS of the Equation 13. Nevertheless, as shown in Supp.Mat. I.2, it wouldn’t be possible to rearrange the $\check{\mathbf{P}}^{(\ell)}$ to recover a logistic function with renormalized parameters. Hence, the resulting method would no longer belong to the logistic parametric family. In other words, LPCA is not self-consistent under coarse-graining.
In addition, note that the calculation of $\check{\mathbf{P}}^{(\ell)}$ requires a greater complexity than the MSM Supp.Mat. subsection II.7.
Secondly, by refitting LPCA at level $\ell$, one would find another set of vectors that, in general, are unrelated with the ($\ell$-1)-vectors. To the point of view of LPCA, the different levels $\ell$ are realizations of different generative process.
To enforce relatedness of the $\ell-$vectors, we forced the lower-resolution parameters $\vec{x}_{I},w_{I}$ to be a function of the higher-resolution ones. Since there is no natural way of combining them, we used the Equation 15, namely
$\vec{b}_{I}:=\sum_{i_{0}\in I}\vec{b}_{i_{0}},\quad\vec{c}_{I}:=\sum_{i_{0}\in I}\vec{c}_{i_{0}}$ (17)
where $I:=\mathbf{\Omega}_{0\to\ell}(i_{0})=(\mathbf{\Omega}_{\ell}\circ\,\cdots\,\circ\mathbf{\Omega}_{0})(i_{0})$ are the block-nodes at level $\ell$. Then, to imposed self-consistency, the parameters are inserted in LPCA activation function, i.e.
$\sigma(\vec{b}_{I},\vec{b}_{J},\vec{c}_{I},\vec{c}_{J})=\frac{1}{1+e^{-\left(\langle\vec{b}_{I},\vec{b}_{J}\rangle-\langle\vec{c}_{I},\vec{c}_{J}\rangle\right)}}$ (18)
At this point, the machinery provides self-consistency and relatedness of LPCA across scales as the MSM does naturally. Hence, we can fairly compare their expected values with respect to $\mathbf{A}^{(\ell\geq 1)}$.
V.3 Comparison Among Probabilities
For every level $\ell$, both LPCA and MSM give rise to 3 probability functions [^6] For the MSM, the symbols on top of $\mathbf{P}$ refers to different values of the inner parameters since its functional form does not vary by construction.: the fitted $\hat{\mathbf{P}}^{(\ell)}$, summed $\widetilde{\mathbf{P}}^{(\ell)}$ and the coarse-grained $\check{\mathbf{P}}^{(\ell)}$. Hence, we produced a cross comparison among them to understand their hallmarks.
Firstly, we will compare for each model how the $\widetilde{\mathbf{P}}^{(\ell)}$ relates with $\check{\mathbf{P}}^{(\ell)}$, i.e. Equation 18 against Equation 8 and Equation 16 against Equation 14.
Secondly, $\widetilde{\mathbf{P}}^{(\ell)}$ against $\hat{\mathbf{P}}^{(\ell)}$, namely Equation 18 against $\hat{\mathbf{P}}^{(\ell)}_{LPCA}$ and Equation 16 against $\hat{\mathbf{P}}^{(\ell)}_{MSM}$.
The insets, displayed in some figure, are reporting the $2D$ histogram of the density of points inside each bin [^7] We set the number of bins equal to $30$ both along the x- and y- axis., i.e.
$r_{xy}=2\frac{n_{\textnormal{bin}_{xy}}}{N_{2}(N_{2}-1)}\in[0,1].$ (19)
where $n_{\textnormal{bin}_{xy}}$ (“xy” refers to its center of mass) is the total number of points and $N_{2}(N_{2}-1)$ the number of pairs. Finally, we colored the bins according to $r_{xy}$ (creating a heatmap) - the bigger the value, the lighter the color.
The missing evaluation of $\hat{\mathbf{P}}^{(\ell)}$ against $\check{\mathbf{P}}^{(\ell)}$ is due to the fact the coarse-grained probability spoils the LPCA functional form (see subsection V.2) whereas $\check{\mathbf{P}}^{(\ell)}=\widetilde{\mathbf{P}}^{(\ell)}$ for the MSM. Therefore, we used it only to check numerically the Equation 13 in the first comparison.
V.4 Scores
V.5 Network Measurements
The fundamental topological properties of a network are the degree, the average nearest neighbor degree (ANND) and the binary clustering coefficient (CC) [8]. Formally, each of such measurements is a function $Y(\mathbf{A})$ of an $N\times N$ adjacency matrix representing a graph $\mathbf{G}$.
Here, we compute them both in the observed network $\mathbf{A}$ and as expected by the model $\mathbf{P}(\mathbf{A}|\textbf{X})$. More precisely, the degree counts the number of edges that are incident to a node $i$, i.e.
$k_{i}(\mathbf{A}):=\sum_{j(\neq i)}a_{ij}$ (20)
and its expected value is given by
$\langle k_{i}\rangle=\sum_{j(\neq i)}p_{ij}$ (21)
[…]
$k^{nn}_{i}(\mathbf{A}):$ $=\sum_{j(\neq i)}\frac{a_{ij}k_{j}}{k_{i}}$ (22)
$=\frac{\sum_{j(\neq i),k(\neq j)}a_{ij}a_{jk}}{\sum_{j(\neq i)}a_{ij}}$ (23)
whereas its expected value reads
$\langle k^{nn}_{i}\rangle$ $:=\langle\frac{\sum_{j(\neq i),k(\neq j)}a_{ij}a_{jk}}{\sum_{j(\neq i)}a_{ij}}\rangle$ (24)
$\approx 1+\frac{\sum_{j(\neq i),k(\neq j,i)}p_{ij}p_{jk}}{\sum_{j(\neq i)}p_{ij}}$ (25)
where in the second passage we took advantage on the first order approximation $\mathbb{E}[\frac{X}{Y}]\approx\frac{\mathbb{E}[X]}{\mathbb{E}[Y]}$ (delta approximation) [30]. Lastly, the CC is defined as the ratio among the number of triangles of node $i$ and its number of wedges, namely
$c_{i}(\mathbf{A})$ $:=\frac{\triangle_{i}}{\wedge_{i}}$ (26)
$=\frac{\sum_{i\neq j\neq k}a_{ij}a_{jk}a_{ki}}{\sum_{j\neq k}a_{ij}a_{ik}}$ (27)
whereas the expected one is
$\langle c_{i}\rangle$ $:=\langle\frac{\triangle_{i}}{\wedge_{i}}\rangle$ (28)
$\approx\frac{\langle\triangle_{i}\rangle}{\langle\wedge_{i}\rangle}$ (29)
$\approx\frac{\sum_{i\neq j\neq k}p_{ij}p_{jk}p_{ki}}{\sum_{j\neq k}p_{ij}p_{ik}}$ (30)
For more complicated measurements, e.g. the variance of the ANND, the delta approximation won’t be valid and one has to estimate them as the average over a sufficiently large ensemble $\mathcal{A}:=\left\{\mathbf{A}_{s}\right\}_{s\in[0,\mathcal{S}-1]}$ where $\mathcal{S}$ is the number of graphs. In particular, having optimized the parameters of the model, we can generate unbiased realizations $\mathcal{A}$ by sampling each $a_{ij}$ independently with probability $p_{ij}$ [30, 17].
In the limit of $\mathcal{S}\to\infty$, the sampled average of any measure $Y_{i}$ meets its analytical estimations $\langle Y_{i}\rangle$ [17], i.e.
$\bar{Y_{i}}:$ $=\frac{1}{\left|\mathcal{A}_{N}\right|}\sum_{\mathbf{\hat{A}}\in\mathcal{A}}Y_{i}(\mathbf{\hat{A}})$ (31)
$\to\langle Y_{i}\rangle$ $=\sum_{\mathbf{B}\in\mathcal{A}_{N}}\mathbf{P}(\mathbf{B}|\textbf{X})Y_{i}(\mathbf{B})$ (32)
where $\mathbf{B}\in\mathcal{A}_{N}$ is a matrix drawn from the set of the undirected binary graphs $\mathcal{A}_{N}$ of $N$ nodes.
Lastly, to estimate the uncertainty of the model over the sampled realizations, we calculated the $97.5$-th ($2.5$-th) percentile of $Y_{i}(\mathcal{A})$ calculated with linear approximation (see [31]). These values are seen as upper and lower bounds of the dispersion intervals $\Delta_{c}(\langle Y_{i}\rangle)$ [26] which contains $c=95\%$ of the measurements $Y_{i}(\mathcal{A}):=\left\{Y_{i}(\mathbf{A}_{s})\right\}_{s\in[0,\mathcal{S}-1]}$ over the sampled graphs. Note that in the whole procedure we arbitrarily fix the percentage of “dispersion” to $c=95\%$ [26], but other values are also allowed.
V.6 Reconstruction Accuracy
In order to have a cross-comparison among all the levels and models, we exploited the reconstruction accuracy [26]. This measure is defined as the fraction of times an observed statistics $Y_{i}$ falls within the dispersion interval $\Delta_{c}(\langle Y_{i}\rangle)$ (see subsection V.5). More formally, the reconstruction accuracy at level $\ell$ for the statistics $Y$ is defined as
$RA^{\ell}_{s}:=\frac{1}{N_{\ell}}\sum_{i=0}^{N_{\ell}-1}\mathbb{I}\left\{Y_{i}\in\Delta(\langle Y_{i}\rangle)\right\}$ (33)
where $\mathbb{I}$ is the indicator function [^8] Further refinements are possible, but we stick with this definition for the sake of simplicity.. Roughly, it counts the frequency at which the sampled ensemble includes the observed statistics. If all the observed statistics, e.g. degrees, were included in the interval, the accuracy would be $1$, whereas the accuracy would be 00 if none of them were included.
V.7 Rescaled ROC and PR Curves
The LPCA and MSM could be seen as binary classifiers that predict the presence of a link between two nodes. For this reason, we evaluated them also for the common metrics used in the Machine Learning field: the expected confusion matrix, the Receiver Operating Characteristic (ROC) and the Precision-Recall (PR) curves [32, 33]. Firstly, the expected confusion matrix is a $2\times 2$ matrix that reports the expected value of True Positives (TP), i.e. $\langle TP\rangle:=\sum_{i\leq j}a_{ij}p_{ij}$, False Positives (FP), i.e. $\langle FP\rangle:=\sum_{i\leq j}(1-a_{ij})p_{ij}$, True Negatives (TN), i.e. $\langle TN\rangle:=\sum_{i\leq j}\left(1-a_{ij}\right)\left((1-p_{ij})\right)$, and False Negatives (FN), i.e. $\langle FN\rangle:=\sum_{i\leq j}a_{ij}\left(1-p_{ij}\right)$ [34]. By combining these scores, one recovers the True Positive Rate (TPR), the False Positive Rate (FPR) and the Positive Predictive Value (PPV) [33], namely
$TPR:$ $=\frac{TP}{Pos}=\frac{\sum_{i<j}a_{ij}p_{ij}}{L}$ (34)
$FPR:$ $=\frac{FP}{Neg}=$ (35)
$=\frac{\sum_{i<j}(1-a_{ij})p_{ij}}{\binom{N}{2}-L}$ (36)
$PPV:$ $=\frac{TP}{PP}\approx\frac{TP}{PP}=\frac{\sum_{i<j}a_{ij}p_{ij}}{\sum_{i<j}p_{ij}}$ (37)
$=\frac{\sum_{i<j}a_{ij}p_{ij}}{L}$ (38)
where $Pos:=\sum_{i<j}a_{ij}=L$, $Neg:=\sum_{i<j}(1-a_{ij})=\binom{N}{2}-L$. It’s possible to “activate” these scores by mapping the entries $p_{ij}\geq\epsilon$ to $1$ and otherwise to 00. For convenience, we will call them $TPR(\epsilon),FPR(\epsilon),PPV(\epsilon)$. The ROC and PR curves are obtained by spanning $\epsilon\in[0,1]$ [32] and plotting, respectively, the TPR against the FPR, and the TPR against the PPV (see Figure 4). Interestingly, if $\epsilon=1$, then $TPR(\epsilon=1)=FPR(\epsilon=1)=0,PPV(\epsilon=1):=1$, whereas if $\epsilon=0$, $TPR(\epsilon=0)=FPR(\epsilon=0)=1,PPV(\epsilon=0)=p:=\frac{Pos}{Pos+Neg}=\frac{L}{\binom{N}{2}}$.
As a reference model, it is commonly employed a random classifier predicting the majority class, i.e.
$argmax_{ij}[p_{ij},1-p_{ij}]$. In the ROC plane, this naive model spans the identity line, and an “L” shape in the PR plane which depends on the link density [^9] As the density increases the corner will be right-shifted and vertical line bent forming a “$\_\backslash_{\_}$” shape. Nonetheless, since we are interested in ranking the summed models, we rescaled the Area Under the Curve (AUC) to highlight the advantage with respect to the naive classifier. Specifically, we defined
$AUC-ROC_{norm}$ $=\frac{AUC_{ROC}-0.5}{0.5}$ (39)
$AUC-PR_{norm}$ $=\frac{AUC_{PR}-p}{1-p}.$ (40)
[…]
V.8 Triangle Density
Inspired by [15], we computed the expected number of triangles for every model at disposal [^10] In this essay, our objective was to model probabilistically the observed network rather than describing it exactly, namely the limit where $p_{ij}\equiv a_{ij}\,\forall i>j$.. Specifically, the expected density of triangles at a certain level $\ell$ is defined as
$\rho^{(\ell)}(c):=\frac{\triangle(\mathbf{G}^{(\ell)}_{k_{i_{\ell}}\geq c})}{2N_{\ell}}=\frac{\sum_{i\neq j\neq k}g_{ij}g_{jk}g_{ki}}{2N_{\ell}}$ (41)
where $\triangle(\mathbf{G}^{(\ell)}_{k_{i_{\ell}}\geq c})$ is the number of observed triangles (see Equation 26) calculated on the subgraph $\mathbf{G}^{(\ell)}_{k_{i_{\ell}}\geq c}$ composed by the nodes $I_{c}:={\left\{i_{\ell}:k_{i_{\ell}}\geq c\right\}}$ with degree lower (or equal) than a threshold $c$ [14, 15]. Its expected value reads
$\langle\rho^{(\ell)}(c)\rangle=\frac{\langle\triangle(\mathbf{G}^{(\ell)}_{k_{i_{\ell}}\geq c})\rangle}{2N_{\ell}}\approx\frac{\sum_{i<j<k}\tilde{p}_{ij}\tilde{p}_{jk}\tilde{p}_{ki}}{N_{\ell}}$ (42)
where $i\in I_{c},j\in I_{c},k\in I_{c}$ and the probabilities $\tilde{p}_{ij}$ refers to the summed model $\widetilde{\mathbf{P}}^{(\ell)}$.
VI Results and Discussions
Here, we present the results of the LPCA and MSM models applied to the ION and WTW datasets.
VI.1 Scale-Invariance Evidence and Multi-Scale Clustering Coefficient
(a) Level 0, 2 - Summed / Renormalized VS Coarse-Grained
(b) Level 0, 2 - Binary Clustering Coefficient
Figure 2: Left: Visualization of the Equation 13 where the LHS lies on the y-axis and the RHS on the x-axis. Right: Cross comparison among LPCA-(8,8) and MSM-16 focusing on the CC. The upper panel reports the expected $\left\{\langle CC_{i}\rangle\right\}_{j\in[1,N]}$ at level 0 whereas the lower at level 2. In addition, the first row refers to the LPCA-(8,8) and the second to MSM-16. The inset highlights the $\widetilde{\mathbf{P}}^{(\ell)}$ against $\hat{\mathbf{P}}^{(\ell)}$ at that resolution level, namely 00 or $2$.
(a) Level 2 - LPCA-(8,8) - Summed, Fitted, Observed Network Measurements
(b) Level 2 - MSM-16 - Summed, Fitted, Observed Network Measurements
Figure 3: Fundamental Network Measurements at level 2 for the ION dataset. In the upper panel, the x-axis hosts the observed measurements whereas the y-axis the expected ones. In particular, from the left one finds the DEG, the ANND and the CC. The single inset depicts the scattered plots of $\widetilde{\mathbf{P}}^{(2)}$ against $\hat{\mathbf{P}}^{(2)}$. In the lower panel, it is displayed the behavior of the network measurements as the degrees increase either for the observed quantities and the expected ones. One finds every expected value calculated with the fitted $\hat{\mathbf{P}}^{(2)}$ (blue) and the summed $\widetilde{\mathbf{P}}^{(2)}$ (orange) while the observed measures are depicted in azure. The left half of the figure refers to the LPCA-(8,8) model, whereas the right half of the figure to the MSM-16.
In Figure 2a, it is represented the behavior of the summed probability against the coarse-grained (see Equation 13) as described in subsection V.3. We chose the lowest embedding dimensions $D_{B}=D_{C}=1$ and $D=1$ to highlight the differences among the models even in the simplest case. Specifically, the identity line depicts the scale-invariant nature (Equation 13) which is met only by MSM. On the other hand, LPCA systematically underestimates the coarse-grained version, since the 0-vectors, maximizing the likelihood at $\ell=0$, get lower values than the ones needed to enforce scale-invariance.
By taking a sufficiently large embedding dimensions $D_{B}=D_{C}=8$ and $D=16$, we reported in Figure 2b a cross-comparison between the LPCA-(8,8) and MSM-16 focusing on the multi-scale CC. At a plain eye inspection, LPCA-(8,8) is outperforming MSM-16 at level 0 (where we fitted the models); whereas it is the other way around at level $2$. More quantitatively at level 00, the $AUC-ROC_{LPCA}\approx 0.92,AUC-PR_{LPCA}\approx 0.89$ while $AUC-ROC_{MSM}\approx 0.91,AUC-PR_{MSM}\approx 0.88$. Additionally, we exploited the relative Froebenius error among the fitted probability $\hat{\mathbf{P}}^{(\ell)}$ and the adjacency matrix defined as [15]
$\epsilon:=\frac{\left\|\hat{\mathbf{P}}-\mathbf{A}\right\|_{2}}{L}$ (43)
to obtain that $\epsilon_{LPCA}\approx 48\%<\epsilon_{MSM}\approx 51\%$. The comparison is further enriched by the insets displaying the summed $\widetilde{\mathbf{P}}^{(\ell)}$ against the fitted $\hat{\mathbf{P}}^{(\ell)}$ at $\ell=0,2$ either for LPCA and MSM - the identity line represents the perfect match. Technically, we created a $2D$ histogram as described in subsection V.3.
Here, we verify numerically that LPCA is not scale-invariant and the agreement of the expected CCs with respect to the observed ones. Comparable results were also obtained for the other levels and measures, namely the DEG and ANND - for ION at $\ell=2$ (see Figure 3), but refer to the Supplementary Material for the WTW and the other levels. Hence, as expected from the theory, the MSM provides the best modelling of the multi-scale ION.
VI.2 Expected Values of the Network Measurements
In Figure 3, we display the key network properties at level $2$ (see subsection V.5) calculated for the empirical network $\mathbf{A}^{(2)}$, as expected by the summed model $\widetilde{\mathbf{P}}^{(2)}$ and by the refitted one $\hat{\mathbf{P}}^{(2)}$ (see section III and V.2). The dimension $D$ we used are $D_{B}=8,D_{C}=8$ for the LPCA (LPCA-(8,8)) and $D=16$ for the MSM (MSM-16).
We computed the observed properties as in subsection V.5. Then, focusing on the fitted model $\hat{\mathbf{P}}^{(\ell=2)}$, we sampled $N_{\mathcal{A}}=1000$ realizations in order to
calculate $\left\{\overline{k_{i_{2}}},\overline{annd_{i_{2}}},\overline{c_{i_{2}}}\right\}_{i_{2}\in[1,N_{2}]}$ as reported in Equation 31. In addition, we obtained the dispersion interval for each measurement with $c=95\%$ the bars attached to every sampled average in the plot. Lastly, we re-applied the same procedure to the summed model $\widetilde{\mathbf{P}}^{(2)}$.
In the upper panel, we report the observed measurements (x-axis) and the expected ones (y-axis) both for the summed model (orange points) and the fitted one (blue points). Needless to say, the identity line represents the perfect match of the predicted quantities with the observed properties. The single inset depicts the scattered plots of $\widetilde{\mathbf{P}}^{(2)}$ against $\hat{\mathbf{P}}^{(2)}$ as described in subsection VI.1.
As expected from Figure 2a, LPCA does not recover the measurements on average whereas the MSM can approximate them including most of the observed points in the dispersion intervals. In addition, by looking at the inset in the upper-left plot, $\widetilde{\mathbf{P}}^{(2)}$ approximates $\hat{\mathbf{P}}^{(2)}$ only for the MSM. This result, differently from Equation 13, was not enforced theoretically, and it explains the good agreement of the MSM measurements through the coarse-grained levels.
In the lower panel, it is displayed the behavior of the network measurements as the degrees increase. From the left-most plot, one finds the complementary cumulative distribution function (CCDF) of the degrees, the ANND and the CC. Since the real CCDF is decreasing, the observed network has a higher presence of lower-degree nodes than the hubs whereas it is not scale-free as its shape is not a straight line in log-log scale. Similarly, the ION (WTW) is disassortative and hierarchical since, respectively, the high-degree nodes are connected to low-degree ones, and they trade with loosely-interacting partners [35].
These plots underscore that the MSM-16, not only capture the CC (see Figure 2b), but also the lower-hops measurements and behaviors. As expected, the LPCA-(8,8) provides a good fit only at the fitted scale as depicted by the blue points for $\hat{\mathbf{P}}^{(0)}$ and $\hat{\mathbf{P}}^{(2)}$.
VI.3 Reconstruction Accuracy and ROC-PR Curves
(a) Level 2 - LPCA-(8,8) - Summed, Fitted ROC and PR curves
(b) Level 2 - MSM-16 - Summed, Fitted ROC and PR curves
Figure 4: Confusion matrices, ROC and PR curves at level 2 for the ION dataset. The left most side is occupied by the two confusion matrices for the fitted $\hat{\mathbf{P}}^{(2)}$ (upper) and the summed $\widetilde{\mathbf{P}}^{(2)}$. The middle plot (ROC curves) reports the behavior of the TPR VS FPR as the threshold goes from 1 to 0 (reading the graphic left-to-right). As before, the two curves are associated with the fitted (blue) model and the summed (orange) one. The left half of the figure refers to the LPCA-(8,8) model, while the right half of the figure to the MSM-16.
(a) All level Reconstruction Accuracy
(b) AUC-ROC and AUC-PR curves through levels
Figure 5: Left: Reconstruction Accuracy (y-axis) by model, level and network statistics. Right: Area Under the ROC and PR curves for the summed models by diminishing the number of nodes or, equivalently, increasing the scale. More concretely, the third point reports the AUC-ROC and PR from Figure 4 as y-coordinates while the x-coordinates are the number of nodes $N_{2}$.
In Figure 5a, one can find the reconstruction accuracy (see Equation 33) for the DEG, ANND, CC across the available levels for the ING network, i.e. $\ell=0,\cdots,3$ [26]. In particular, we reported the summed LPCA with dimensions $D_{B}=D_{C}=1$ and $D_{B}=D_{C}=8$, and the summed MSM with $D=1,2,8,16$. Since we have fitted every model at $\ell=0$, only most of the trends are peaked at the resolution scale, for e.g. the MSM-1 has a higher CC at level $1$ than at 00. In addition, out of the fitted level, the LPCA fails in generating an ensemble of networks that are consistent with the observed one. Contrarily, the MSM ensemble includes the measures at every level apart from resolution $2$ where the topology change could not be grasped only by the 0-parameters and the renormalization rule. Especially for $D=1$, the MSM overestimates the DEG, ANND, CC since the summation of the 0-parameters lead to bigger values than the 2-parameters fitted at level $2$. As said, this is not per se a problem of the MSM since the scale-invariance enforces the inner consistency of the MSM (see Equation 13) rather than recovering the fitted parameters at every level. Also, the choice of the pathological partition leads to worse results than expected as discussed in subsubsection VI.3.1. In addition, by increasing the number of parameters, i.e. $D$, the reconstruction accuracy improves, but it leads to overfitting as highlighted by the higher BIC scores reported in Supp.Mat. Table 1. Another way to tackle this deviation, could be to introduce a dyadic relationship $d_{ij}$ among the nodes as done in [17]. However, this is out of the scope of this work.
[…]
In conclusion, the MSM can consistently model all the coarse-grained levels, whereas the LPCA outperforms the MSM only at the fitting scale $\ell=0$. This implies that, by fine tuning the functional form, one can prioritize either the single-scale “overfitting” with LPCA or the generalization capability with the MSM.
VI.3.1 Dependence on an arbitrary partition
The agreement of the summed model and the observed (coarser) graph depends on the chosen partitions. In other words, by recalling the notation used in as in Equation 1 and Equation 13, the relationship
$a_{IJ}$ $\approx p_{IJ}$ (44)
iff (45)
$1-\prod_{i_{0}\in I,j_{0}\in J}(1-a_{i_{0}j_{0}})$ $\approx 1-\prod_{i_{0}\in I,j_{0}\in J}\left(1-p\left(x_{i_{0}},x_{j_{0}},w_{i_{0}}\right)\right)$ (46)
depends on the partition since $I,J$ are functions of $\Omega$. Therefore, one can engineer a partition $\Omega^{diff}$ that spoils the latter approximation by requiring that $p_{IJ}$ addresses the zeros of $a_{IJ}$ (cfr. Equation 44) through
$\Omega^{diff}:=\min_{\Omega}\sum_{I\leq J}\left(\prod_{i_{0}\in I,j_{0}\in J}(1-p_{i_{0}j_{0}})-a_{IJ}\right)^{2}$ (47)
As said, $\Omega^{diff}$ would provide a partition leading to worse results than the ones observed for ION and WTW.
VI.4 Expected Number of Triangles
(a) Level 0
(b) Level 2
Figure 6: Left: Rich-Club-Triangle Density at level 0. The plot shows the evolution of the $\rho^{(\ell)}(c)$ with respect to the degree $k$ in azure solid dots. The other markers identify the expected $\langle\rho^{(\ell)}(c)\rangle$ by the models: LPCA-(1,1) (azure circles), LPCA-(8,8) (orange triangles), MSM-1 (green circles), MSM-2 (red circles), MSM-8 (violet crosses), MSM-16 (brown triangles). Right: Rich-Club-Triangle Density at level 2.
The authors of [15] show that it is possible, by introducing LPCA, to reproduce the triangle density (TriDens) for an embedding dimension lower that the number of nodes (cfr. [14]). Here, we show the expected TriDens as described by Equation 41, Equation 42.
In Figure 6, the filled azure dots depict the observed TriDens whereas the other markers identify each model: azure circles (LPCA-(1,1)), orange triangles (LPCA-(8,8)), green circles (MSM-1), red circles (MSM-2), violet crosses (MSM-8), brown triangles (MSM-16). In Figure 6a, it is reported the level $\ell=0$ where it is clear that even the lowest embedding dimension ($D=1$) well approximates the TriDens. By construction, as $c$ increases, the difference among the TriDens vanishes until it coincides at $c=N-1$ since the subgraph of hubs is likely not to contain triangles in a disassortative network (see Figure 3). This result is not in contrast with the previous works [14, 15] since we are computing the expected TriDens rather than the exact one. By coarse-graining the network Figure 6b at $\ell=2$, the MSM models are in a good agreement with the coarser TriDens. In particular, as seen in Figure 3, as $D$ increases, also the estimates improve. On the contrary, LPCA are underestimating the considered score being biased because, as said previously, it is not generalizable to lower resolutions.
VII Conclusions
The power of graphs relies on their ability to accommodate different kind of interactions by a suitable definition of nodes and edges. By arbitrarily identifying a node as the aggregation of microscopic entities the resulting network is a coarse-grained version of the original one. By repeating this procedure several times, one obtains a multi-scale unfolding of the observed graph. We have applied this procedure on the ION and the WTW (see subsection IV.1) showing how one generative process could be represented at several resolutions.
A relevant part of the node-embedding literature [36] aims at find the optimal representation of the nodes at a single scale, neglecting these numerous ways of tracking down the generative process. To stress the point, we applied either a single-scale method from the machine-learning field, i.e. LPCA [16], and the new multi-scale models enriched with node embeddings (MSM) section III. The key assumption is that, similarly to the generative process, the models must be scale-invariant through the scales. Since the MSM is built with this premise, it would naturally be self-consistent whereas we forced LPCA to be self-consistent by applying the same renormalization rule of the MSM (Equation 17). In particular, the renormalization rule (Equation 15) states that the community vectors are the sum of their inner-node vectors. This allows for a principled interpretation of the sum of the node embeddings which is not possible with LPCA and, in general, with the single-scale models.
At fitting scale $\ell=0$, LPCA outperforms MSM in every metric we have considered in section V. At higher scales, the ranking is reversed (see Figure 5) as the predictions of LPCA highly deviates from the observed structure - this defined this change as single-scale overfitting. More specifically, in Figure 2a, we visualized at which extent imposing the self-consistency makes LPCA diverge from its coarse-graining probability - the MSM satisfies this identically. Secondly, we showed the agreement (disagreement) of the expected network measurement by using the summed MSM (LPCA). This implies that LPCA has to be fitted at every level as if every scale would be generated by a different generative process. Therefore, also the node embeddings would not be related to each other. On the contrary, the MSM can be fitted at the highest resolution, providing the fundamental vectors that can be summed to obtain the higher-level embeddings. This is a clear advantage of the MSM over the LPCA - even computationally as reported in the Supp.Mat. subsection II.7.
As cross-comparison between the models, we visualized the Reconstruction Accuracy (Figure 5a), the AUCs (Figure 5b) and the expected number of triangles (Figure 6). Specifically, the Reconstruction Accuracy highlights that the ensembles, generated by the two models, includes the observed quantities at level $\ell=0$. However, by changing scale, LPCA generates graphs that are not related to the empirical one. Interestingly, also the MSM struggles to recover the observed properties at level $\ell=2$ because the graph topology changes more than what expected by the model. In particular, without the dyadic parameters, the MSM overestimates the 0-parameters which lead to a higher density of edges at level $2$ (see the z-score in the legend of Figure 3b). As a consequence, all the measurements are overestimated and the dispersion interval can’t include the observed values. In Figure 5b, it has been depicted the clear sign of single-scale overfitting: the predictability of LPCA is restricted to $\ell=0$ since the AUCs constantly decrease when it is applied to higher levels. Lastly, the analysis of the expected triangular density (Figure 6) shows that it is possible to generate networks with comparable values as the observed triangular density. As said before, this possibility is spoiled under aggregation for the LPCA, but it is preserved for the MSM.
In conclusion, LPCA, developed using the maximum Shannon entropy principle, is the best model at its fitted level but struggles to accurately represent the network’s structure at other scales. In contrast, the multi-scale model (MSM), based on the scale-invariance principle, consistently captures coarser resolutions, such as the ION and WTW. The decline in LPCA’s predictive performance at higher scales suggests that MSM provides a better balance for modeling the multi-scale structures. In addition, the MSM offers a meaningful interpretation of node embedding sums, as they naturally generate lower-resolution levels, making it a more versatile and comprehensive approach for analyzing networks.
VII.1 Future Perspectives
Although this work used undirected and binary models, it was a good starting point to extend the analysis to directed methods [37] and, hence, the interpretation of the resulting directed node embeddings within the economic theory. For the weighted part, there is still theoretical work to do to understand how the weights could be included in the MSM framework.
VIII ACKNOWLEDGMENTS
We thank ING Bank N.V. for their support and active collaboration. A special thanks to the whole DataScience team at ING Bank for their advice that helped shape this research.
SUPPLEMENTARY MATERIAL
accompanying the paper
“Multi-Scale Node Embeddings for Graph Modeling and Generation”
by R. Milocco, F. Jansen and D. Garlaschelli
I Non-Negative Logistic PCA
The non-negative LPCA (LPCA) [15] aims to classify every edge $(i,j)$ as existing (0) or non-existing (1). By treating every entry $a_{ij}$ of the adjacency matrix A as a Bernoulli random variable (Equation 5), this comes down to the factorization
$\textbf{A}\sim\sigma(\textbf{B}\textbf{B}^{T}-\textbf{C}\textbf{C}^{T})\qquad\textnormal{or}\qquad a_{ij}\sim\sigma(\langle\vec{b}_{i},\vec{b}_{j}\rangle-\langle\vec{c}_{i},\vec{c}_{j}\rangle):=\frac{1}{1+e^{-\left(\langle\vec{b}_{i},\vec{b}_{j}\rangle-\langle\vec{c}_{i},\vec{c}_{j}\rangle\right)}}\quad\forall i,j$ (1)
where $\sigma$ is the logistic function depending on the scalar product of two embedding per nodes assumed to encode the role of each node in the network, namely $\vec{b}_{i}\in\mathbb{R}_{+}^{D_{B}},\vec{c}_{i}\in\mathbb{R}_{+}^{D_{C}}\textnormal{ where }D_{B}\geq 0,D_{C}\geq 0$. The compact formulation on the LHS was written in terms of the matrices $\textbf{B}\in\mathbb{R}_{+}^{N\times D_{B}},\textbf{C}\in\mathbb{R}_{+}^{N\times D_{C}}$ that are created by stacking horizontally the vectors $\vec{b}_{i}\textnormal{ and }\vec{c}_{i}$ respectively.
Similarly, for the MSM (section III), the LPCA vectors are fitted by means by maximizing the log-likelihood estimation” [38]. In particular, the log-likelihood and its gradient read
$\mathcal{L}\left(\mathbf{B},\mathbf{C}|\mathbf{A}\right):=$ $\sum_{i\leq j}a_{ij}\ln\left(\sigma_{ij}\right)+\left(1-a_{ij}\right)\ln\left(1-\sigma_{ij}\right)$ (2)
$=$ $\sum_{i\leq j}\min(x_{ij},0)-\ln(1+e^{-\left|x_{ij}\right|})-(1-a_{ij})x_{ij}$ (3)
$\partial_{b_{ik}}\mathcal{L}=$ $\sum_{j}\left(a_{ij}-\sigma_{ij}\right)b_{jk}$ (4)
$\partial_{c_{ik}}\mathcal{L}=$ $-\sum_{j}\left(a_{ij}-\sigma_{ij}\right)c_{jk}$ (5)
where the last passage is taken from BCE-TensorFlow for numerical stability. Note that the stationarity conditions
$\partial_{b_{ik}}\mathcal{L}\stackrel{{\scriptstyle!}}{{=}}0$ (6)
$\partial_{c_{ik}}\mathcal{L}\stackrel{{\scriptstyle!}}{{=}}0$ (7)
[…]
I.1 Renormalizing the LPCA
The LPCA does not have a recipe to renormalize the parameters and produce an “up-scaled” version of it. Nonetheless, it is possible to model the multi-level structure $\left\{\mathbf{A}_{\ell}:\ell\geq 0\right\}$ either by fitting $\hat{\mathbf{P}}_{\ell}$ at every level or by coarse-graining ($\check{\mathbf{P}}^{(\ell)}$) as given by
$\check{p}^{(\ell)}_{IJ}:=1-\prod_{i_{0}\in I,j_{0}\in J}\frac{1}{1+e^{\langle\vec{b}_{i_{0}},\vec{b}_{j_{0}}\rangle-\langle\vec{c}_{i_{0}},\vec{c}_{j_{0}}\rangle}}$ (8)
where $I:=\mathbf{\Omega}_{0\to\ell}(i_{0}),J:=\mathbf{\Omega}_{0\to\ell}(j_{0})$ are the block-nodes at level $\ell$. However, $\mathbf{P}_{cg}^{(\ell)}$ spoils the self-consistency [^11] As said, $\left\{\mathbf{A}_{\ell}:\ell\geq 0\right\}$ provides multiple representation of the same generative process. Hence, the model should mimic this feature with self-consistency. of LPCA; a property that allows one functional form for all the levels. In particular, Equation 8 is a product of logistic functions which is not re-writtable as a logistic function.
Since we want to test the capability of the model to remain self-consistent, we will renormalize the parameters by summing the microscopic parameters Equation 17 as done for the MSM Equation 15 and check for its agreement with the empirical quantities.
As pointed out in subsection II.7, in order to compute $\check{\mathbf{P}}^{(\ell)}$, the computational complexity is higher than the summed $\widetilde{\mathbf{P}}^{(\ell)}$. Therefore, for the fairest comparison among the models, we will use $\widetilde{\mathbf{P}}^{(\ell)}_{LPCA}$.
I.2 Inconsistency of the LPCA: a trivial example
By referring to Figure 1, the microscopic network of 3 nodes $0,1,2$ merges into the community $A,B$ containing respectively the nodes $0,1$ and the node $2$, namely $A=\Omega(0)=\Omega(1),B=\Omega(2)$.
00level 0 11112222
$\qquad\stackrel{{\scriptstyle\textnormal{Coarse-Graining}}}{{\longrightarrow}}\qquad$ ABlevel 1
Figure 1: Trivial Network for Inconsistency of LPCA and CM. The blue nodes belong to community $A$ whereas the red node to the community $B$.
In addition, to stress the point we will switch-off the dependence on $\vec{c}_{i}$. Therefore, the connection probability of the communities $A,B$ from the Equation 18 as
$\sigma(b_{AB})=\left(1+e^{-b_{AB}}\right)^{-1}$
which is different from the coarse-grained (Equation 8)
$\sigma^{cg}(b_{AB})$ $=1-\frac{1}{1+e^{-b_{0}b_{2}}}\frac{1}{1+e^{-b_{1}b_{2}}}$
$=\frac{e^{-b_{0}b_{2}}+e^{-b_{1}b_{2}}+e^{-b_{0}b_{2}}e^{-b_{1}b_{2}}}{1+e^{-b_{0}b_{2}}+e^{-b_{1}b_{2}}+e^{-b_{0}b_{2}}e^{-b_{1}b_{2}}}.$
From the previous results, it is clear that from $\sigma^{cg}(b_{AB})$ one can’t recover the $\sigma(b_{AB})$ by defining $b_{A}:=f(b_{0},b_{1}),b_{B}:=b_{2}$ similarly to Equation 15. Therefore, the model is not renormalizable.
II Derivation of the multi-scale probability
Here, we derive the multi-scale model formulation enhanced with vectors (MSM). As in the main text, we will consider a coarse-graining procedure from 00 to $\ell\geq 0$ even though the treatment will hold for every pair $m,\ell+1$ with $m\leq\ell$. See [17], for further details even for the following passages.
Before introducing the model, it is worth to recall the problem settings to generate the observed multi-scale structure. Concretely, let us consider the binary undirected adjacency matrix $\mathbf{A}^{(0)}$ at level 00 describing the microscopic interactions among the 0-nodes. Subsequently, a hierarchical and non-overlapping partition of the microscopic nodes $\left\{\mathbf{\Omega}_{\ell}\right\}_{\ell\geq 0}$ prescribing the community (block-nodes) membership of the lower-level nodes. Concretely, the block-nodes $I:=i_{\ell}$ hosting all the $i_{0}$ nodes is obtained by
$I:=\mathbf{\Omega}_{0\to\ell-1}(i_{0})$
where we have defined $\Omega_{0\to\ell-1}:=\Omega_{\ell-1}\circ\dots\circ\Omega_{0}$. Lastly, a rule to assign a link among blocks, namely
$a_{IJ}=1-\prod_{i_{0}\in I,j_{0}\in J}(1-a_{i_{0}j_{0}})$ (9)
where $a_{IJ}:=a^{(\ell)}_{i_{\ell+1}j_{\ell+1}},a_{i_{0}j_{0}}:=a^{(0)}_{i_{0}j_{0}}$ and $J:=j_{\ell}=\mathbf{\Omega}_{0\to\ell-1}(j_{0})$. Therefore, by iterating the procedure it is possible to create the nested set of networks describing the original phenomenon at different resolutions.
In order to model this architecture, one needs several assumptions. The first one requires that the MSM must describe the microscopic matrix $\textbf{A}^{(0)}$. Similarly to section I, $\textbf{A}^{(0)}\sim P^{(0)}\big{(}\mathbf{A}^{(0)},\mathcal{X}^{(0)}\big{)}$ subject to $P^{(0)}\big{(}\cdot,\mathcal{X}^{(0)}\big{)}=(P^{(0)})^{T}\big{(}\cdot,\mathcal{X}^{(0)}\big{)}$ and $\sum_{\textbf{S}\in\left|\mathscr{A}^{(0)}\right|}P^{(0)}\big{(}\textbf{S},\mathcal{X}^{(0)}\big{)}=1$ where $\mathscr{A}^{(0)}$ is the ensemble of all the binary symmetric graphs with $N_{0}$ nodes.
In line with [17], we further assume that
$\mathcal{X}^{(0)}_{ij}:=\begin{cases}\langle\vec{x}_{i},\vec{x}_{j}\rangle&\textnormal{if }i\neq j\\[8.61108pt] \frac{1}{2}\left\|\vec{x}_{i}\right\|^{2}+w_{i}&\textnormal{if }i=j\end{cases}$ (10)
is given by a product (to-be-defined) between the $D-$dimensional vectors $\left\{\vec{x}_{i}\right\}_{i\in[1,N_{\ell}]}$ with additional node-wise parameters $\left\{w_{i}\right\}_{i\in[1,N_{\ell}]}$ only active in the self-loop part ($i_{\ell}=j_{\ell}$). In particular, $\vec{x}_{i}$ encodes the capability of node $i$ to connect to the other nodes; whereas $w_{i}$ its propensity for a self-interaction. Since the principles leading to an edge are different to the self-loop ones, we introduced two independent parameters. Furthermore, the matrix of parameters $\mathcal{X}^{(0)}$ could include also a dyadic relationship among the nodes. The higher order terms have been discarded since they will not be fully compatible with the hypothesis of “independent edges”. For further details we refer to [17].
To highlight the parameter dependence of the model, in the following, we will use the notation $P^{(0)}\big{(}\mathbf{A}^{(0)},\mathcal{X}^{(0)}\big{)}:=P^{(0)}\big{(}\mathbf{A}^{(0)},\mathbf{X}_{w}^{(0)}\big{)}$ where $\mathbf{X}^{(0)}_{w}:=[\textbf{X},\vec{w}]$. Technically, $\mathbf{X}^{(0)}:=\left[\vec{x}_{1},\dots,\vec{x}_{N_{0}}\right]^{T}\in N_{0}\times D$ and $\vec{w}^{(0)}:=\left\{w_{i_{0}}\right\}_{i\in[1,N_{0}]}$.
Fitted $\mathbf{X}^{(0)}_{w}$ at the ground level, the ensemble generated by the MSM contains multiple configurations that, after coarse-graining, lead to the observed macroscopic $\mathbf{A}^{(\ell)}$ [17], i.e. $\{\mathbf{A}^{(0)}\}\xrightarrow{\mathbf{\Omega_{0\to\ell-1}}}\mathbf{A}^{(\ell)}$. In turns, this induces the probability of observing $\mathbf{A}^{(\ell)}$ as
$P_{\ell}\big{(}\mathbf{A}^{(\ell)},\mathbf{X}^{(0)}_{w}\big{)}:=\!\!\!\!\!\!\!\!\!\!\!\!\!\sum_{\{\mathbf{A}^{(0)}\}\xrightarrow{\Omega_{0\to\ell-1}}\mathbf{A}^{(\ell)}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!P_{0}\big{(}\mathbf{A}^{(0)},\mathbf{X}^{(0)}_{w}\big{)}.$ (11)
To enforce the scale-invariance property, we require that the functional form of the MSM has to be independent from the chosen scale, i.e. $P_{\ell}\big{(}\cdot,\cdot)\stackrel{{\scriptstyle!}}{{=}}P_{0}\big{(}\cdot,\cdot)\quad\forall\ell\geq 0$. Furthermore, that the model can generate the possible $\ell$-graphs in two equivalent ways hierarchically or directly. The former one refers to Equation 11, and it prescribes to generate the 00-graph ensemble with probability $P\big{(}\mathbf{A}^{(0)},\mathbf{X}^{(0)}_{w}\big{)}$ and, then, coarse-graining them $\ell$ times via the partitions $\{\mathbf{\Omega}_{k}\}^{\ell-1}_{k=0}$. The other way around, the second one requires to renormalize the parameters $\widetilde{\mathbf{X}}_{\widetilde{w}}^{(\ell)}$ and, then, directly model $\mathbf{A}^{(\ell)}$ via $P\big{(}\mathbf{A}^{(\ell)},\widetilde{\mathbf{X}}_{\widetilde{w}}^{(\ell)})$. Imposing both requirements
$P\big{(}\mathbf{A}^{(\ell)},\widetilde{\mathbf{X}}_{\widetilde{w}}^{(\ell)}\big{)}$ =!$\stackrel{{\scriptstyle!}}{{=}}\!\!\!\!\!\!\!\!\!\!\!\!\!\sum_{\{\mathbf{A}^{(0)}\}\xrightarrow{\mathbf{\Omega}_{0\to\ell-1}}\mathbf{A}^{(\ell)}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!P\big{(}\mathbf{A}^{(0)},\mathbf{X}_{w}^{(0)}\big{)}\quad\Leftrightarrow\quad\widetilde{\mathbf{P}}^{(\ell)}\stackrel{{\scriptstyle!}}{{=}}\check{\mathbf{P}}^{(\ell)}$ (12)
where we have defined the LHS and RHS of the first equation as $\widetilde{\mathbf{P}}^{(\ell)}\textnormal{ and }\check{\mathbf{P}}^{(\ell)}$ respectively. In other words, the form of the $P(\cdot,\cdot)$ will depend on the scale $\ell$ only through the renormalized parameters $\widetilde{\mathbf{X}}_{\widetilde{w}}^{(\ell)}$. Moreover, by assuming that the links are statistically independent, the previous equation yields
$p_{IJ}$ =$\stackrel{{\scriptstyle!}}{{=}}\check{p}_{IJ}$ (13)
where $p_{IJ}:=p_{IJ}(\widetilde{\mathbf{X}}_{\widetilde{w}}^{(\ell)})$ and
$\check{p}_{IJ}:=1-\prod_{i_{0}\in I,j_{0}\in J}(1-p\left(\vec{x}_{i_{0}},\vec{x}_{j_{0}},w_{i_{0}}\right))$ (14)
depend, respectively, on the renormalized and the fitted parameters at level $\ell=0$. Note that we didn’t use $\widetilde{p}_{IJ}$ because the functional form would be scale-invariant and the only dependence on the scale is through the parameters $IJ$. The interpretation is similar to the Equation 9: the probability $p_{IJ}$ that there is one among the block-nodes $I,J$ is given by the probability that there is at least one link among the microscopic nodes $i_{0}\in I,j_{0}\in J$. Specifically, it requires that the model remains self-similar whereas the parameters renormalize under renormalization (scale-variant).
The RHS returns the coarse-grained probability for every model, e.g. the MSM and LPCA. The crucial difference is that for the SSM $\widetilde{\mathbf{P}}^{(\ell)}\neq\check{\mathbf{P}}^{(\ell)}$ (cfr. Equation 14) because they are scale-invariant. For a concrete example, we refer to the sections where the models have been introduced.
By taking the logarithm of both sides of the Equation 13, the only functional form compatible with that constraint
$\ln(1-p_{IJ})=-\langle g(\vec{x}_{I}),g(\vec{x}_{J})\rangle$
[…]
$\ln(1-p_{IJ})=-\langle\vec{x}_{I},\vec{x}_{J}\rangle$
one ends up with the (off-diagonal) scale-invariant probability Equation 16. For the self-loops part, the steps are similar to the ones described in the main reference.
II.1 Bilinearity Requirement
00level 0 111122223333
$\qquad\stackrel{{\scriptstyle\textnormal{Coarse-Graining}}}{{\longrightarrow}}\qquad$ ABlevel 1
Figure 2: Simple graph where the nodes $0,1$ merges into the community $A$ whereas $2,3$ in the block-node $B$. The gray dashed lines represent the non-existing links.
In this subsection, we describe why the $\langle$*,$\rangle$ ∗ , ∗ ⟩ product must be bilinear and why M can be taken as the identity matrix given that the probability is symmetric. To start with, in Figure 2, there have been represented 4 nodes $0,1,2,3$ at level $\ell=0$ merging, at level $\ell=1$, into $A:=\left\{0,1\right\}$ and $B:=\left\{2,3\right\}$. Hence, from Equation 13, the non-existence of a link (gray dashed lines) among the communities $A\textnormal{ and }B$ reads
$e^{-\langle\vec{x}_{A},\vec{x}_{B}\rangle}\stackrel{{\scriptstyle!}}{{=}}\prod_{i\in A;j\in B}e^{-\langle\vec{x}_{i},\vec{x}_{j}\rangle}\quad\textnormal{ iff }e^{-\langle\vec{x}_{1}+\vec{x}_{1},\vec{x}_{2}+\vec{x}_{3}\rangle}\stackrel{{\scriptstyle!}}{{=}}e^{-\langle\vec{x}_{1},\vec{x}_{2}\rangle+\langle\vec{x}_{1},\vec{x}_{3}\rangle+\langle\vec{x}_{1},\vec{x}_{2}\rangle+\langle\vec{x}_{1},\vec{x}_{3}\rangle}$ (15)
and the rightmost side enforces that the $\langle$*,$\rangle$ ∗ , ∗ ⟩ has to be a bilinear function.
As said in the main text, the connection probability is symmetric, namely $\mathbf{P}\stackrel{{\scriptstyle!}}{{=}}\mathbf{P}^{T}$. In turns, this leads to $\mathbf{M}=\mathbf{M}^{T}$. In addition, since $p_{ij}\in[0,1]\,\forall i,j$, M is also positive semidefined as $\vec{x}_{i}\textbf{M}\vec{x}_{j}\geq 0\;\forall i,j$. Therefore, $\mathbf{M}$ has positive eigenvalues such that
$\vec{x}_{i}^{T}\textbf{M}\vec{x}_{j}=\vec{x}_{i}^{T}\mathbf{O}\mathbf{O}^{T}\vec{x}_{j}=\vec{y}_{i}^{T}\vec{y}_{j}$
where $\vec{y}_{i}:=\mathbf{O}^{T}\vec{x}_{i}\in\mathbb{R}_{+}^{D}$ (cfr. Cholesky decomposition). Briefly, choosing an arbitrary (symmetric) $\mathbf{M}$ matrix will lead to $\vec{x}_{i}=\textbf{O}\vec{y}_{i}$ with $\vec{y}_{i}$ are optimized with $M:=\textnormal{Id}_{D\times D}$. Hence, for simplicity, we rely on $\mathbf{M}:=\textnormal{Id}_{D\times D}$.
Lastly, fixing $\mathbf{M}:=\textnormal{Id}_{D\times D}$, allows recovering the product among scalars $x_{i}x_{j}$ for $D=1$. This was a successful way of modelling real world networks, e.g. [17, 30].
II.2 Loop parameters estimate
The log-likelihood regarding the self-loops reads
$\mathcal{L}\left(\mathbf{X}_{w}|\mathbf{A}^{diagonal}\right)=\sum_{\left\{i\textnormal{ s.t. }a_{ii}=1\right\}}\ln\left(1-e^{-\frac{1}{2}\|\vec{x}_{i}\|^{2}-w_{i}}\right)-\sum_{\left\{i\textnormal{ s.t. }a_{ii}=0\right\}}\left(\frac{1}{2}\|\vec{x}_{i}\|^{2}+w_{i}\right)$ (16)
which depends either on vectors $\left\{\vec{x}_{i}\right\}_{i\in[1,N_{\ell}]}$ but also on $w_{i}$. Since the probability is bounded, $w_{i}\geq-\frac{1}{2}\|\vec{x}_{i}\|^{2}$. Moreover, as for every node $i$ there would be only one of the two terms in the above likelihood, $w_{i}$ is going to take the values reported in Equation 13. Hence, after having obtained the $\left\{\vec{x}_{i}\right\}_{i\in[1,N_{\ell}]}$, we fixed $w_{i}$ to exactly reproduce the self-loops at level $\ell=0$ as prescribed by Equation 13.
II.3 Gradient of the log-likelihood
To efficiently calculate the maximum of the MSM likelihood, one needs the analytical expression of the gradient. In particular, by differentiating with respect to the $t$-th component of the $s$-th embedding vector, one gets
$\partial_{st}\mathcal{L}$ $=\sum_{i\leq j}\left(\frac{a_{ij}}{p_{ij}}-1\right)\partial_{st}\langle\vec{x}_{i},\vec{x}_{j}\rangle$ (17)
$=\sum_{i\leq j}\left(\frac{a_{ij}}{p_{ij}}-1\right)\left(x_{it}\delta_{sj}+x_{jt}\delta_{si}\right)$ (18)
$=\frac{1}{2}\sum_{i\neq j}\left(\frac{a_{ij}}{p_{ij}}-1\right)\left(x_{jt}\delta_{si}+x_{it}\delta_{sj}\right)+\sum_{i}\left(\frac{a_{ii}}{p_{ii}}-1\right)x_{it}\delta_{si}$ (19)
$=\sum_{j}\left(\frac{a_{sj}}{p_{sj}}-1\right)x_{jt}$ (20)
where we have used
$q_{ij}$ $:=e^{-\langle\vec{x}_{i},\vec{x}_{j}\rangle}$ (22)
$\frac{\partial p_{ij}}{\partial x_{st}}$ $=q_{ij}\left[x_{jt}\delta_{is}+x_{it}\delta_{js}\right]$ (23)
$\frac{\partial\ln q_{ij}}{\partial x_{st}}$ $=-\frac{\partial(x_{i}x_{j})}{\partial x_{st}}=-\frac{1}{q_{ij}}\frac{\partial p_{ij}}{\partial x_{st}}.$ (24)
Thus, by renaming the indexes,
$\partial_{ik}\mathcal{L}=\sum_{j}\left(\frac{a_{ij}}{p_{ij}}-1\right)x_{jk}\equiv\sum_{j}\left(\frac{a_{ij}}{1-e^{-\langle\vec{x}_{i},\vec{x}_{j}\rangle}}-1\right)x_{jk}$ (25)
Lastly, by leveraging on the gradient and the likelihood, we performed the optimization by means of three optimizers: Adam implemented from [39]; while Truncated-Conjugate Gradient and L-BFGS-B by means of the SciPy library [40].
II.4 Structural Equivalence is not Statistical Equivalence for multidimensional-node embeddings
Calculating the gradient Equation 25 for nodes $i\textnormal{ and }i^{\prime}$ at the maximum of the likelihood, one gets
$\partial_{ik}\mathcal{L}\stackrel{{\scriptstyle!}}{{=}}0\stackrel{{\scriptstyle!}}{{=}}\partial_{i^{\prime}k}\mathcal{L}.$ (26)
By further assuming $i\textnormal{ and }i^{\prime}$ have same neighbors (Structural Equivalence - StructE), i.e. $\mathcal{N}(i)=\mathcal{N}(i^{\prime})$, it is possible to rearrange the above equations into
$\phi(x_{ik}):=\sum_{j\in\mathcal{N}(i)}\frac{x_{jk}}{p_{ij}}=\sum_{j\in\mathcal{N}(i)}x_{jk}=\sum_{j\in\mathcal{N}(i^{\prime})}\frac{x_{jk}}{p_{i^{\prime}j}}=:\phi(x_{i^{\prime}k}).$ (27)
where we have defined
$\phi(y_{ik})=\sum_{j\in\mathcal{N}(i)}\frac{y_{jk}}{1-e^{-\langle\vec{y}_{i},\vec{x}_{j}\rangle}}.$ (28)
For $D=1$,
$\phi(x_{i}):=\sum_{j\in\mathcal{N}(i)}\frac{x_{j}}{1-e^{-x_{i}x_{j}}}$ (29)
is a monotonic function of $x_{i}$, thus there exists an inverse function $\phi^{-1}$ such that $\phi(x_{i})\stackrel{{\scriptstyle!}}{{=}}\phi(x_{i^{\prime}})\Rightarrow x_{i}=\phi^{-1}(\phi(x_{i}))=x_{i}$. In other words, two nodes $i,i^{\prime}$ that have the same neighbors, they are statistically equivalent (StatE). The inverse implication is also true, i.e. $x_{i}=x_{i^{\prime}}\Rightarrow\mathcal{N}(i)=\mathcal{N}(i^{\prime})$ by following the same steps backwards. In conclusion, for $D=1$, Structural Equivalence is equivalent to Statistical Equivalence: $x_{i}=x_{i^{\prime}}\textnormal{ if, and only if }\mathcal{N}(i)=\mathcal{N}(i^{\prime})$.
This result does not hold for $D>1$ since $\phi(y_{i}k)$ is not monotonic in $y_{i}k$ as it depends on the scalar product $\langle\vec{y}_{i},\vec{x}_{j}\rangle$. However, if $i,i^{\prime}$ have the same neighbors, they have the same role in the network, and so the model should have the same parameters for the two nodes. Formally,
$\mathcal{N}(i)=\mathcal{N}(i^{\prime})\Leftrightarrow a_{ij}\equiv a_{i^{\prime}j}\Rightarrow p_{ij}=p_{i^{\prime}j}\Leftrightarrow\vec{x}_{i}=\vec{x}_{i^{\prime}}\qquad\forall j\in[1,N_{\ell}].$
By recalling the notation used before, we are imposing that StructE implies StatE.
To find the node embeddings in the reduce problem, one starts by defining the set of nodes which share the same neighbors as $\mathcal{S}:=\left\{\mathcal{S}_{i},\,\forall i\in[0,N-1]\right\}$ where $\mathcal{S}_{i}:=\left\{j:\mathcal{N}(j)=\mathcal{N}(i)\right\}$ are the nodes with the same neighborhood with respect to the node $i$. Referring to Figure 2 and considering the gray dashed lines as existing connection, $\mathcal{S}_{0}:={0,1}$ and $\mathcal{S}_{2}:={2,3}$. Secondly, by defining the node $i$ with the lowest index among $S_{i}$, as the representative of that StructE class, one obtains a set of pivotal nodes related to StructE. At this point, the number of parameters decreased from $N\times D$ to $N_{\mathcal{S}}\times D$ where $N_{\mathcal{S}}$ is the number of elements the set $\mathcal{S}$ has. In order to recover $\mathbf{X}_{w}^{(\ell)}$, the value of the representative parameter $\vec{x}_{i}$ is copied to all the members of the class, formally
$\left\{\vec{x}_{i}:=\vec{x}_{i_{r}}\,\forall i:\mathcal{N}(i)\equiv\mathcal{N}(i_{r})\right\}\qquad\forall i_{r}\in[0,N_{S}-1]$
where $i_{r}$ are the representative nodes. In this way, it is possible to use the likelihood Equation 11 with fewer parameters than in the original formulation.
Since the selected optimizers ([40]) are not “stochastic”, one may obtain the StatE between the embeddings by starting from the same initial conditions for StructE nodes. Indeed, the parameter updates will be the same as they depend on the neighborhood as shown by Equation 4, 25.
In conclusion, for $D=1$, the StructE - StatE relied on the monotonic function $\phi(x_{k})$; whereas it is not the case for $D>1$. Therefore, to enforce StatE, one may reduce the problem and optimize only the representative of the StructE classes.
II.5 Removal of deterministic nodes
[…]
II.6 From Constrains to Bounds
The MSM probability requires a positive inner products
$\langle\vec{x}_{i},\vec{x}_{j}\rangle\geq 0$ (30)
for every pair of nodes, in order to guarantee $p_{ij}\in[0,1]\forall i,j$.
Here, we will prove that the above constraints is equivalent of setting all vector components to be non-negative, namely $x_{ik}\geq 0$. Roughly, the spanned region by the embeddings is enclosed in one quadrant of the space, and it is possible to rotate the vectors to lay in the positive quadrant.
The steps to show this are the following. First, since we are interested on the sign among the vectors, one may restrict to the set of unit vectors $\left\{\vec{e}_{j}\right\}_{i\in[0,N-1]}$ where $\vec{e}_{i}:=\frac{\vec{x}_{i}}{\left\|\vec{x}_{i}\right\|}\in\mathbb{R}^{D}$ and assume $\vec{e}_{0}:=[1,0,\dots]$. Taking into consideration also Equation 30, the considered set is
$\mathbb{S}:=\left\{\vec{e}_{i}\in\mathbb{R}^{D}\textnormal{ s.t. }0\leq\langle\vec{e}_{i},\vec{e}_{j}\rangle\leq 1,\vec{e}_{0}:=[1,0,\dots]\quad\forall i\in[0,N-1]\textnormal{ and }j\in[0,N-1]\right\}$
which by construction has to property that
$\max_{i\in[1,N_{\ell}],j\in[1,N_{\ell}]}\left|\arccos(\langle\vec{e}_{i},\vec{e}_{j}\rangle)\right|\leq\frac{\pi}{2}$
If one takes the most “clockwise” and “anticlockwise” vectors in the set $\mathbb{S}$ defined as
$\vec{e}_{c}$ :=$:=\left\{\vec{e}_{i}:\vec{e}_{i}\times v\geq 0,\,\forall v\in\mathbb{S}\right\}$ (31)
$\vec{e}_{ac}$ :=$:=\left\{\vec{e}_{i}:\vec{e}_{i}\times v\leq 0,\,\forall v\in\mathbb{S}\right\},$ (32)
by construction they form an angle $\theta_{a-ac}\in[-\frac{\pi}{2},\frac{\pi}{2}]$. Therefore, if $\vec{e}_{c}=\vec{e}_{0}$ the treatment is finished since the vectors lay in the positive quadrant. On the other hand, if $\vec{e}_{ac}=\vec{e}_{0}$, the vectors lay in the negative quadrant and they can be rigidly rotated to lie on the positive quadrant, i.e. $x_{ik}\geq 0$.
II.7 Algorithmic Complexity
As described in the subsection I.2, in order to describe a coarser graph without refitting the parameters, one should to use the RHS of Equation 13121212This spoils its functional form, but this would be the only way to avoid refitting the same model at a higher scale. Specifically, the algorithmic complexity to obtain one $\check{p}_{IJ}$ is $cN_{I}N_{J}$ where the evaluation of $p_{ij}$ is assumed to have a complexity $c$. Hence, to compute the complexity of $\check{\mathbf{P}}^{(\ell)}$, one has to sum over all the pairs, i.e.
$\mathcal{C}_{SSM}=c\sum_{I>J}N_{I}N_{J}=\frac{c}{2}\sum_{I,J}N_{I}N_{J}=c\frac{N^{2}_{0}}{2}$ (33)
where $N_{0}$ are the number of structural inequivalent nodes at $\ell=0$. On the other hand, assuming the “sum” of vectors in Equation 15 of order $O(1)$, the complexity of $\widetilde{\mathbf{P}}^{(\ell)}$ (see Equation 16) reads
$\mathcal{C}_{MSM}\approx N_{0}(D+1)+c\frac{N_{\ell}(N_{\ell}-1)}{2}$ (34)
where the summation over the $w_{i}$ parameters counts as $N_{0}$ operations and $\binom{N_{\ell}}{2}$ are all pairs at the level $\ell$ for the full $\widetilde{\mathbf{P}}^{(\ell)}$.
The improvement is reported by the ratio of the two complexities, namely
$\frac{\mathcal{C}_{MSM}}{\mathcal{C}_{SSM}}\approx\frac{2(D+1)}{cN_{0}}+\frac{N_{\ell}\left(N_{\ell}-1\right)}{N_{0}^{2}}\approx\frac{1}{N_{0}}\left[1+\frac{N_{\ell}\left(N_{\ell}-1\right)}{N_{0}}\right].$ (35)
Focusing on level $\ell=3,N_{3}=87,N_{0}=972$,
$\frac{\mathcal{C}_{MSM}}{\mathcal{C}_{SSM}}\approx\mathcal{O}(10^{-2})$ (36)
one saves two order of magnitude by proceeding with the summed MSM rather than the coarse-grained LPCA. Hence, by means of Equation 17, LPCA recovers the same complexity of summed MSM; thereby enabling for a comparison of equal complexity.
II.8 Principled Embedding Dimension via Information Criteria
(a) AICs by model class for ION Model Dim Level 0 Level 1 Level 2 Level 3 LPCA (1, 1) 0.5764 0.6661 0.5448 0.3733 (8, 8) 0.5072 0.5779 0.4025 2.6648e+17 MSM 1 0.5784 0.6658 0.5408 0.4097 2 0.5516 0.6341 0.5129 0.3841 3 0.5434 0.6242 0.5012 0.3791 4 0.5378 0.6169 0.4984 0.3831 5 0.5336 0.6127 0.4985 0.4046 6 0.5311 0.6088 0.4979 0.4232 7 0.5304 0.6074 0.4968 0.4574 8 0.5287 0.6052 0.4976 0.5118 9 0.53 0.6073 0.498 0.5588 10 0.5305 0.6064 0.5002 0.6091 11 0.5311 0.6105 0.5117 0.6671 16 0.5365 0.6226 0.5328 0.9697
(a) BICs by model class for ION Model Dim Level 0 Level 1 Level 2 Level 3 LPCA (1, 1) 0.622 0.7295 0.6605 0.7189 (8, 8) 0.8718 1.0856 1.3276 2.6648e+17 MSM 1 0.6012 0.6975 0.5987 0.5825 2 0.5972 0.6976 0.6285 0.7296 3 0.6117 0.7194 0.6746 0.8974 4 0.6289 0.7439 0.7297 1.0741 5 0.6476 0.7714 0.7877 1.2684 6 0.6678 0.7992 0.8449 1.4597 7 0.69 0.8295 0.9015 1.6667 8 0.7111 0.8591 0.9602 1.8939 9 0.7352 0.8929 1.0184 2.1136 10 0.7584 0.9237 1.0785 2.3367 11 0.7818 0.9595 1.1477 2.5675 16 0.9011 1.1303 1.4579 3.7339
Table 1: Normalized AIC and BIC scores tables for the models LPCA and MSM at different levels for the ION. The best scores are highlighted in green whereas the worst in red.
(a) AICs by model class for WTW
Model Dim Level 0 Level 1 Level 2 Level 3 Level 4 Level 5
LPCA (1, 1) 0.6109 0.5636 0.572 0.5764 0.5584 0.444
(8, 8) 0.4536 0.4309 0.5289 0.7033 1.0492 2.0645
MSM 1 0.6146 0.5637 0.5656 0.5649 0.5387 0.3991
2 0.5807 0.5332 0.5389 0.5285 0.5016 0.3665
3 0.5684 0.5483 0.5528 0.5519 0.5446 0.4753
4 0.5567 0.5366 0.5522 0.5642 0.5645 0.4876
5 0.5628 0.5295 0.5511 0.567 0.5825 0.5279
6 0.5674 0.5356 0.5603 0.5946 0.6059 0.607
7 0.5707 0.5424 0.5679 0.6174 0.6557 0.7085
8 0.5181 0.4733 0.4959 0.5363 0.5764 0.7742
9 0.5809 0.5649 0.6051 0.6526 0.7167 0.8852
10 0.5923 0.5765 0.6198 0.6819 0.7609 0.994
11 0.6017 0.5949 0.6362 0.7114 0.7962 1.0688
16 0.557 0.5571 0.631 0.7292 0.9881 1.5484
(b) BICs by model class for WTW
Model Dim Level 0 Level 1 Level 2 Level 3 Level 4 Level 5
LPCA (1, 1) 0.7813 0.7583 0.8004 0.8551 0.922 0.9868
(8, 8) 1.8166 1.9881 2.3555 2.9326 3.958 6.4068
MSM 1 0.6974 0.6578 0.677 0.6982 0.7088 0.6026
2 0.7463 0.7215 0.7616 0.795 0.8417 0.7736
3 0.8169 0.8306 0.8869 0.9518 1.0548 1.0859
4 0.8881 0.9131 0.9976 1.0972 1.2448 1.3018
5 0.977 1.0001 1.1079 1.2334 1.4329 1.5456
6 1.0645 1.1003 1.2284 1.3942 1.6264 1.8282
7 1.1506 1.2013 1.3473 1.5503 1.8462 2.1333
8 1.1809 1.2263 1.3867 1.6025 1.9369 2.4025
9 1.3265 1.4121 1.6073 1.8521 2.2474 2.7171
10 1.4208 1.5177 1.7333 2.0147 2.4616 3.0295
11 1.513 1.6303 1.8611 2.1774 2.6669 3.3078
16 1.8825 2.0631 2.4126 2.8616 3.7093 4.8051
Table 2: Normalized AIC and BIC scores tables for the models LPCA and MSM at different levels for the WTW. The best scores are highlighted in green whereas the worst in red.
To determine the “best” embedding dimension for LPCA and MSM, we used the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC) [27], [28]. These scores are defined as
AIC $:=2K-2\ln(\mathcal{L})$ (37)
BIC $:=K\ln(n)-2\ln(\mathcal{L})$ (38)
where $K$ is the number of parameters, $n:=\binom{N}{2}$ the number of observations and $\mathcal{L}$ the likelihood of the model. They encode the trade-off between the goodness-of-fit $-\ln(\mathcal{L})$ and the complexity of the model. Therefore, the “best model” is the one with the minimum AIC or BIC; which one of the two remains a debated choice: the AIC is asymptotically equivalent to the Kullback-Leibler divergence among the generating model and a candidate one [27] whereas the BIC to the Description Length (DL) [28]. As the DL is the only one embodying the “trade-off” paradigm, we decided to select the minimum BIC criterion. In Equation 38, the scores are not comparable across scale, therefore, we normalized the AIC and BIC scores as
$\textnormal{AIC}_{\textnormal{norm},\ell}$ $:=\frac{2}{n_{\ell}}K_{\ell}-\frac{2\ln(\mathcal{L})}{n_{\ell}}$ (39)
$=4\left[\frac{D}{N_{\ell}-1}-\frac{\ln(\mathcal{L})}{N_{\ell}\left(N_{\ell}-1\right)}\right]$ (40)
$\textnormal{BIC}_{\textnormal{norm},\ell}$ $:=\frac{\ln(n_{\ell})}{n_{\ell}}K_{\ell}-\frac{2\ln(\mathcal{L})}{n_{\ell}}$ (41)
$\approx\frac{\ln(N_{\ell})+\ln(N_{\ell}-1)}{N_{\ell}-1}D-4\frac{\ln(\mathcal{L})}{N_{\ell}(N_{\ell}-1)}$ (42)
where $K_{\ell}=N_{\ell}D$ and $n_{\ell}:=\binom{N_{\ell}}{2}$. This doesn’t affect the ranking, but it provides the AIC and BIC per pair.
The results are summarized in Table 1 for the ION and Table 2 for the WTW where the best scores (minimum) are highlighted in green whereas the worst (maximum) in red. The comparison is provided among LPCA and MSM as they have a different functional forms especially in the combination of the parameters. We have considered only two levels for LPCA as the benchmark for lowest $D$ and “maximum” $D$ with respect to our computational facility. On the other hand, we spanned more dimensions for MSM to provide an extensive description of the model performances. Recall that, by increasing the level of coarse-graining, the network tends to be less complex: the Equation 9 likely densifies the network implying that the nodes will have more similar roles. Therefore, the ideal dimension $D$ decreases. Lastly, the average BIC score is calculated to provide a global view of the model performances. Thus, the best model is the one with the lowest average BIC score across levels. Overall, the best models are LPCA-(1,1) and MSM-1 as BIC penalizes more than AIC the complexity of the model. However, in the following we will display the behavior also of the $D=2,8,16$ for MSM and $D=(8,8)$ to assess the model performances at higher dimensions.
II.9 Statistical Impossibility into the application of Train and Test Split
Since every pair is seen as a Bernoulli random variable not identically distributed [41, 8], this implies that every link has to be accounted in the training procedure since it is not already modelled by other pairs.
III World Trade Web results
In this section, we report the same results presented in the main text, but for the World Trade Web. The conclusions are the same as for the ING network.
(a) Level 0, 4 - Summed / Renormalized VS Coarse-Grained
(b) Level 0, 4 - Binary Clustering Coefficient
Figure 3: Left: Visualization of the Equation 13 where the LHS lies on the y-axis and the RHS on the x-axis. Right: Cross comparison among LPCA-(8,8) and MSM-16 focusing on the CC. The upper panel reports the expected $\left\{\langle CC_{i}\rangle\right\}_{j\in[1,N]}$ at level 0 whereas the lower at level 4. In addition, the first row refers to the LPCA-(8,8) and the second to MSM-16. The inset highlights the $\widetilde{\mathbf{P}}^{(\ell)}$ against $\hat{\mathbf{P}}^{(\ell)}$ at that resolution level, namely 00 or $4$.
(a) Level 4 - LPCA-(8,8) - Summed, Fitted, Observed Network Measurements
(b) Level 4 - MSM-16 - Summed, Fitted, Observed Network Measurements
Figure 4: Fundamental Network Measurements at level 4 for the ION dataset. In the upper panel, the x-axis hosts the observed measurements whereas the y-axis the expected ones. In particular, from the left one finds the DEG, the ANND and the CC. The single inset depicts the scattered plots of $\widetilde{\mathbf{P}}^{(4)}$ against $\hat{\mathbf{P}}^{(4)}$. In the lower panel, it is displayed the behavior of the network measurements as the degrees increase either for the observed quantities and the expected ones. One finds every expected value calculated with the fitted $\hat{\mathbf{P}}^{(4)}$ (blue) and the summed $\widetilde{\mathbf{P}}^{(4)}$ (orange) while the observed measures are depicted in azure. The left half of the figure refers to the LPCA-(8,8) model, whereas the right half of the figure to the MSM-16.
(a) Level 4 - LPCA-(8,8) - Summed, Fitted ROC and PR curves
(b) Level 4 - MSM-16 - Summed, Fitted ROC and PR curves
Figure 5: Confusion matrices, ROC and PR curves at level 4 for the ION dataset. The left most side is occupied by the two confusion matrices for the fitted $\hat{\mathbf{P}}^{(4)}$ (upper) and the summed $\widetilde{\mathbf{P}}^{(4)}$. The middle plot (ROC curves) reports the behavior of the TPR VS FPR as the threshold goes from 1 to 0 (reading the graphic left-to-right). As before, the two curves are associated with the fitted (blue) model and the summed (orange) one. The left half of the figure refers to the LPCA-(8,8) model, while the right half of the figure to the MSM-16.
(a) All level Reconstruction Accuracy
(b) AUC-ROC and AUC-PR curves through levels
Figure 6: Left: Reconstruction Accuracy (y-axis) by model, level and network statistics. Right: Area Under the ROC and PR curves for the summed models by diminishing the number of nodes or, equivalently, increasing the scale. More concretely, the third point reports the AUC-ROC and PR from Figure 5 as y-coordinates while the x-coordinates are the number of nodes $N_{4}$.
(a) Level 0
(b) Level 4
Figure 7: Left: Rich-Club-Triangle Density at level 0. The plot shows the evolution of the $\rho^{(\ell)}(c)$ with respect to the degree $k$ in azure solid dots. The other markers identify the expected $\langle\rho^{(\ell)}(c)\rangle$ by the models: LPCA-(1,1) (azure circles), LPCA-(8,8) (orange triangles), MSM-1 (green circles), MSM-4 (red circles), MSM-8 (violet crosses), MSM-16 (brown triangles). Right: Rich-Club-Triangle Density at level 4.